### v1

# On the connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity

Abstract
In this work, we establish a connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity. We begin by defining a convolution product between maps from the contracted semigroup algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ to an arbitrary associative algebra $\mathcal{A}$. The convolution product makes the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to $\mathcal{A}$ an associative algebra. We prove that the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ and the tensor product algebra $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ are isomorphic. As a consequence, in the specific case of the inverse semigroup of matrix units, we identify the product in the space of maps on the matrix algebras which is preserved by the Choi-Jamiołkowski isomorphism as convolution. Then, by defining the Fourier transform of a map from $\mathbb{C}_{0}[S]$ to $M_{n}(\mathbb{C})$, we derive the Fourier inversion formula when $S$ is a finite inverse semigroup. As a corollary of this formula, we show that in the case of the inverse semigroup of matrix units, the Fourier transformation of a map with respect to the identity representation becomes the Choi matrix of the map and the Fourier inversion formula becomes the Choi inversion formula. Then, by defining the notion of matrix valued positive definite maps, we prove Bochner’s theorem in the context of finite inverse semigroup. It is demonstrated that Bochner’s theorem reduces to Choi theorem on completely positive maps when the inverse semigroup of matrix units is considered. Additionally, the necessary and sufficient condition on a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$ is obtained.
I Introduction
In the context of quantum theory, linear maps between spaces of operators are essential for describing the evolution of quantum states. Specifically, a quantum state is represented by a density operator—a positive semidefinite matrix with unit trace. A physical process is then modeled by a linear map that sends density operators to other density operators. However, not every such map corresponds to a valid physical transformation. A key requirement is that the map must be completely positive and trace-preserving. Such maps are called quantum channels. Complete positivity ensures that even when the system is entangled with another (possibly unknown) system, the overall transformation remains valid. This is crucial because in quantum mechanics, systems are often not isolated. A foundational result that provides a practical criterion for complete positivity is the Choi theorem [1]. It states that a linear map is completely positive if and only if its associated Choi matrix is positive semidefinite. This correspondence makes it easier to verify whether a given map can represent a physical process.
The Fourier transform is a fundamental tool in both physics and mathematics, with widespread applications across numerous domains. The Fourier transform of an integrable function $f(x)$ on the real line is defined as: $\widehat{f}(t):=\int_{-\infty}^{\infty}e^{-ixt}f(x)dx$ for all $t\in\mathbb{R}$ with the inverse transform given by $f(x):=\frac{1}{2\pi}\int_{-\infty}^{\infty}e^{ixt}\widehat{f}(t)dt$ for all $x\in\mathbb{R}$. These definitions can be generalized for complex valued function on a finite group as follows: Let $G$ be a finite group and $f:G\to\mathbb{C}$ be a complex valued function. Then the Fourier transform of $f$ with respect to a representation $\rho:G\to M_{d_{\rho}}(\mathbb{C})$ is defined as $\widehat{f}(\rho):=\sum_{g\in G}\rho(g)f(g)$ with the inverse transform given by $f(g)=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}(\rho(g^{-1})\widehat{f}(\rho))$, where $\mathrm{Irr}(G)$ denotes a complete set of inequivalent irreducible representations of $G$. A function $f:G\to\mathbb{C}$ is called positive definite if $\sum_{g,t\in G}f(g^{-1}t)c_{t}\overline{c}_{g}\geq 0$ for any set of complex numbers $\{c_{g}\}_{g\in G}$. Positive definite functions are completely characterized by Bochner’s theorem which states that a function $f:G\to\mathbb{C}$ is positive definite if and only if its Fourier transforms $\widehat{f}(\rho))$ are positive semidefinite for $\rho\in\mathrm{IrrU}(G)$, where $\mathrm{IrrU}(G)$ represents a complete set of inequivalent irreducible unitary representations of $G$. For a map $\Phi:G\to M_{n}(\mathbb{C})$, its Fourier transform with respect to a representation $\rho$ can be defined as $\widehat{\Phi}(\rho):=\sum_{g\in G}\rho(g)\otimes\Phi(g)$. A map $\Phi:G\to M_{n}(\mathbb{C})$ is called positive definite if the matrix $\Phi(g^{-1}g^{\prime})]$ is positive semidefinite. In this case Bochner’s theorem states that a map $\Phi:G\to M_{n}(\mathbb{C})$ is positive definite if and only if its Fourier transforms $\widehat{\Phi}(\rho))$ are positive semidefinite for $\rho\in\mathrm{IrrU}(G)$ (details are provided in the main text).
Now it can be observed that the Choi matrix of a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$, defined as $C_{\Phi}:=\sum_{i,j}e_{ij}\otimes\Phi(e_{ij})$, where $\{e_{ij}\}_{i,j=1}^{m}$ denotes the standard matrix units of $M_{m}(\mathbb{C})$, can be interpreted as the Fourier transform of $\Phi$. In particular, the expression $C_{\Phi}:=\sum_{i,j}\mathrm{id}_{m}(e_{ij})\otimes\Phi(e_{ij})$ closely resembles the standard definition of the Fourier transform of maps on finite groups, as discussed above. Moreover, the convolution-like product between two linear maps $\Phi,\Phi^{\prime}:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$, defined by $(\Phi*\Phi^{\prime})(e_{ij}):=\sum_{k}\Phi(e_{ik})\Phi^{\prime}(e_{kj})$, is preserved under the Choi–Jamiołkowski isomorphism; that is, $C_{\Phi*\Phi^{\prime}}=C_{\Phi}C_{\Phi^{\prime}}$ [2]. This equation resembles the well-known convolution theorem, which states that the Fourier transform of a convolution is the product of the individual Fourier transforms. These structural parallels motivate us to define a Fourier transform for matrix-valued maps on the contracted algebra, prove Bochner’s theorem in the context of finite inverse semigroups, and establish a connection with Choi theorem on completely positive maps.
I.1 Our Contributions and Overview of our Main Results
With the aim to establish a connection between Bochner’s theorem and Choi theorem, we begin by defining a convolution product between maps from the contracted semigroup algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ to an arbitrary associative algebra $\mathcal{A}$. The convolution product makes the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to $\mathcal{A}$ an associative algebra. Then we proved that the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ and the tensor product algebra $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ are isomorphic (see Theorem (1). As a consequence of Theorem (1), in the specific case of matrix unit semigroup, we have identified that the product in the space of maps on the matrix algebras which is preserved by the Choi-Jamiołkowski isomorphism is convolution. Then defining the Fourier transform of a map from $\mathbb{C}_{0}[S]$ to $M_{n}(\mathbb{C})$, we have derived the Fourier inversion formula when $S$ is a finite inverse semigroup in Theorem (2) which is one of our main results. As a corollary of Theorem (2) we have shown that in the case of the inverse semigroup of matrix units, the Fourier transformation of a map with respect to the identity representation becomes the Choi matrix of the map and the Fourier inversion formula becomes the Choi inversion formula [see Corollary (1)]. Additionally, We also have derived the Plancherel formula in the context of inverse semigroup [see Theorem (3)].
Then, by defining the notion of matrix valued positive definite maps, we prove Bochner’s theorem in the context of finite inverse semigroup in Theorem (4) which is one of our main results. While proving Bochner’s theorem for finite inverse semigroup, we have provided two useful alternative characterization of positive definite maps on contracted semigroup algebras in Proposition (3) and Proposition (4). Also, we have proved Schur’s orthogonality relation in the context of finite inverse semigroup in Proposition (5) which played a crucial role in the proof of Bochner’s theorem. Then in subsection (III.4) we have shown how Bochner’s theorem reduces to Choi theorem on completely positive maps. Finally, in Theorem (6), we have derived the necessary and sufficient condition on a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$.
I.2 Structure of the paper
The structure of the paper is as follows: In Section II, we provide basic ideas of a semigroup, inverse semigroup and their contracted algebras. Section III contains our main results. It is divided into several subsections. In Subsection III.1, we define and discuss various properties of Fourier transform of maps on the contracted semigroup algebra of a finite semigroup. In Subsection III.2, we derive the Fourier inversion formula for finite inverse semigroup and show that in the case of matrix unit semigroup, the formula reduces to the Choi inversion formula. Additionally, Plancherel formula in the context of finite inverse semigroup is also derived. By defining the notion of matrix valued positive definite maps, we prove Bochner’s theorem in the context of finite inverse semigroup in Subsection III.3. Also we provide two useful alternative characterizations of positive definite maps on the contracted semigroup algebra of a finite inverse semigroup. Here, we prove Schur’s orthogonality relation in the context of finite inverse semigroup which played a crucial role in the proof of Bochner’s theorem. Then, in subsection III.4, we show how Bochner’s theorem reduces to Choi theorem on completely positive maps. In Subsection III.5, we derive the necessary and sufficient condition on a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$. Finally, we conclude our work in Section IV.
II Preliminaries
In this section, we briefly discuss some basics of finite inverse semigroup, its useful properties and representations. Also we discuss Choi theorem on completely positive maps briefly.
II.1 Semigroup, Inverse Semigroup and their Algebras
A semigroup is a pair $(S,\cdot)$ of a non-empty set $S$ and a binary operation $\cdot:S\times S\to S$ which is associative, i.e., $x\cdot(y\cdot z)=(x\cdot y)\cdot z$ for all $x,y,z\in S$. Throughout the paper, we adopt the notation that $x\cdot y=xy$.
Definition.
(Semigroup Algebra) Let $S$ be a semigroup. A semigroup-algebra of $S$ over the complex field $\mathbb{C}$ is an algebra $\mathbb{C}[S]$ with a basis which is isomorphic to the semigroup $S$ under the multiplication rule in $\mathbb{C}[S]$. With the abuse of notation, we will use $S$ itself to denote the basis for $\mathbb{C}[S]$.
Definition.
A contracted algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ is an algebra over $\mathbb{C}$, which have a basis $\mathscr{B}$ such that $\mathscr{B}\cup\{0\}$ is isomorphic to $S$, where 0 represents the zero of the contracted algebra.
With the abuse of notation, we will use $S\setminus\{0\}$ (or sometimes $S\setminus z$) to denote a basis for $\mathbb{C}_{0}[S]$. In the following, we provide a brief description of the basic concepts and properties of an inverse semigroup. The concept of an inverse semigroup provides the notion of inverse without having the notion of an identity element unlike a group.
Definition A.
(Inverse Semigroup) A semigroup $S$ is called an inverse semigroup if for every element $s\in S$ there exists a unique element denoted by $s^{-1}$ such that $ss^{-1}s=s$ and $s^{-1}ss^{-1}=s^{-1}$.
It trivially follows from the definition that $(s^{-1})^{-1}=s$. An element $p\in S$ is called an idempotent element if $p^{2}=p$. Now we list some elementary properties that hold in an inverse semigroup.
Proposition.
Let $S$ be an inverse semigroup. Then the following properties holds.
  1. 1.
    For any element $s\in S$, the elements $ss^{-1}$ and $s^{-1}s$ are idempotents.
  2. 2.
    If $e\in S$ is an idempotent, then $e^{-1}=e$.
  3. 3.
    Let $e$ and $f$ be two idempotents in $S$. Then $ef$ is also an idempotent and $ef=fe$. i.e., the idempotents commutes.
  4. 4.
    For $s,t\in S$, we have $(st)^{-1}=t^{-1}s^{-1}$.
II.1.1 Partial ordering and Möbius Inversion
Let $S$ be a finite partially ordered set and $R$ be a commutative unital ring. The incidence algebra denoted by $R[S\times S]$ is an algebra of all the functions $f:S\times S\to R$ such that whenever $x\in S$ and $y\in S$ are not compatible i.e., they are not related by the partial ordering in $S$ ($x\nleq y$ or $y\nleq x$), we have $f(x,y)=0$. The multiplication between $f,h\in R[S\times S]$ is the convolution defined as follows:
$f*h(x,y):=\sum_{\begin{subarray}{c}z\in S\\ x\leq z\leq y\end{subarray}}f(x,z)h(z,y).$ (1)
The incidence algebra is a unital algebra with the identity element defined as $\mathrm{I}(x,y)=\delta_{xy}$, where $\delta_{xy}$ is the usual Kronecker delta. With the identity being present in the incidence algebra, the inverse of $f\in R[S\times S]$ can be defined if there exist $f^{-1}\in R[S\times S]$ such $f*f^{-1}(x,y)=f^{-1}*f(x,y)=\mathrm{I}(x,y)$ for all $x,y\in S$ with $x\leq y$. Equivalently
$\sum_{\begin{subarray}{c}z\in S\\ x\leq z\leq y\end{subarray}}f(x,z)f^{-1}(z,y)=\delta_{x,y}.$ (2)
The above equation is used to find $f^{-1}(x,y)$ inductively. Specifically, for all $x\in S$, we have $f^{-1}(x,x)=\frac{1}{f(x,x)}$ from the above equation. Then this data is used to find $f^{-1}(x,z)$ for $z\in S$ such that there doesn’t exist $z_{0}\in S$ except $x$ and $z$ with $x\leq z_{0}\leq z$. Continuing in this way, it is possible to find $f^{-1}(x,y)$.
Definition.
(Zeta function and Möbius function) The zeta function is an element $\zeta$ in $R[S\times S]$ such that $\zeta(x,y)=1$ for all $x\leq y$ and 0 otherwise. The Möbius function denoted by $\mu$ is the inverse of the zeta function.
It is clear that $\mu(x,x)=\frac{1}{\zeta(x,x)}=1$ and when $x<y$, from Eq. (2), we have
$\sum_{\begin{subarray}{c}z\in S\\ x\leq z\leq y\end{subarray}}\zeta(x,z)\mu(z,y)=0,$
$\iff$ $\sum_{\begin{subarray}{c}z\in S\\ x\leq z\leq y\end{subarray}}\mu(z,y)=0,$
$\iff$ $\mu(x,y)=-\sum_{\begin{subarray}{c}z\in S\\ x<z\leq y\end{subarray}}\mu(z,y).$ (3)
The above equation can be used inductively to determine the Möbius function.
Theorem B.
(Möbius Inversion) Let $S$ be a partially ordered set with the ordering denoted by $\leq$ and let $R$ be a field. Let $f:S\to R$ and $h:S\to R$ be two functions satisfying for all $x\in S$
$f(x)=\sum_{\begin{subarray}{c}y\in S\\ y\leq x\end{subarray}}h(y).$ (4)
Then $h$ can be expressed in terms of $f$ as follows:
$h(x)=\sum_{\begin{subarray}{c}y\in S\\ y\leq x\end{subarray}}\mu(y,x)f(y).$ (5)
II.1.2 Steinberg’s Isomorphism
An example of an inverse semigroup is the symmetric inverse semigroup. Denoted by $\mathcal{J}_{X}$, it is the semigroup consists of all partial bijections $\tau:Y\to Z$ on a non-empty set $X$, where $X,Y\subseteq X$. The binary operation is the composition of partial transformations. It can be verified that the $\tau:Y\to Z$ and $\tau^{-1}:Z\to Y$ satisfies all the defining properties of an abstract inverse semigroup defined in Definition (A). Now, the Preston-Vagner representation theorem [3] says that any inverse semigroup $S$ can be thought of as an inverse subsemigroup of $\mathcal{J}_{X}$.
Preston-Vagner representation: Let $S$ be an inverse semigroup and $\mathcal{J}_{S}$ be the symmetric inverse semigroup of all partial bijection of $S$. Then $S$ is isomorphic to an inverse subsemigroup of $\mathcal{J}_{S}$.
For $\tau\in\mathcal{J}_{S}$, we denote by $\mathrm{dom}(\tau)$ the domain of $\tau$ and by $\mathrm{ran}(\tau)$ the range of $\tau$. Denoting by $\mathrm{id}_{U}$ the identity map on the subset $U\subseteq S$, we have $\tau\tau^{-1}=\mathrm{id}_{\mathrm{ran}(\tau)}$ and $\tau^{-1}\tau=\mathrm{id}_{\mathrm{dom}(\tau)}$. As $\mathrm{id}_{U}$ can be identified with the subset $U\subseteq S$ itself, it a customary to define for an abstract inverse semigroup $S$ and $s\in S$
$\mathrm{dom}(s):=s^{-1}s,$ (6)
$\mathrm{ran}(s):=ss^{-1}$ (7)
and think of $s$ as an isomorphism $s:s^{-1}s\to ss^{-1}$. Recalling that $s^{-1}s$ and $ss^{-1}$ are idempotents, the notion of two idempotents being isomorphic is defined as follows:
Definition C.
(Isomorphic idempotents)[4] Let $S$ be an inverse semigroup. The idempotents $e,e^{\prime}\in S$ are said to be isomorphic if there exists $s\in S$ such that $e=s^{-1}s$ and $e^{\prime}=ss^{-1}$.
A partial ordering on an inverse semigroup $S$ is defined as follows.
Definition D.
Let $S$ be an inverse semigroup and $s,t\in S$. Then $s\leq t$ if there exist an idempotent $e\in S$ such that $s=et$.
For two arbitrary elements $s$ and $t$ belonging to an inverse semigroup $S$, the following are equivalent.
  1. 1.
    $s\leq t$.
  2. 2.
    There exists and idempotent $e^{\prime}$ such that $s=te^{\prime}$.
  3. 3.
    $s=ss^{-1}t$.
  4. 4.
    $s=ts^{-1}s$.
Definition E.
[4] Let $S$ be a finite inverse semigroup. For every non-zero element $s\in S$, let us define the following:
$\lfloor s\rfloor:=\sum_{t\in S,t\leq s}\mu(t,s)t\in\mathbb{C}_{0}[S],$ (8)
where $\mu(t,s)$ is the Möbius function defined on the partial order of $S$ as given in Eq. (3).
It simply follows that for the zero element $z\in S$, $\lfloor z\rfloor=0$. It is proved in Ref. [4], that $\{\lfloor s\rfloor\ :s\neq z,s\in S\}$ is a basis for the contracted algebra $\mathbb{C}_{0}[S]$. The basis $\{\lfloor s\rfloor\}_{s\in S\setminus z}$ is known as the groupoid basis. Using the Möbius inversion formula, we can recover the natural basis $\{s\}_{s\in S\setminus z}$ as :
$s=\sum_{t\in S,t\leq s}\lfloor t\rfloor.$ (9)
The basis elements in $\{\lfloor s\rfloor\}_{s\in S\setminus z}$ multiplies as [4]:
$\lfloor s\rfloor\lfloor t\rfloor=\begin{cases}\lfloor st\rfloor&\text{if}\ s^{-1}s=tt^{-1},\\ 0&\text{otherwise}.\end{cases}$ (10)
Definition.
Let $S$ be a finite inverse semigroup. Then, $s,t\in S$, are said to be $\mathcal{D}$-relatable and denote by $s\mathcal{D}t$ if there exist an element $x\in S$ such that $x^{-1}x=ss^{-1}$ and $xx^{-1}=tt^{-1}$
The above defined relation is called Green’s $\mathcal{D}$-relation, and it can be shown that it is an equivalence relation on $S$. The equivalence classes with respect to this relation are known as the $\mathcal{D}$-classes.
Let $\{D_{k}\}_{k=1}^{n}$ be the $\mathcal{D}$-classes of $S$. Let us define $\mathbb{C}D_{k}:=\mathrm{Span}\{\lfloor s\rfloor\}$ where $s\in D_{k}$. It should be noted that there always exists an idempotent inside each $\mathcal{D}$-classes $D_{k}$, because if $t\in D_{k}$ then $t^{-1}t\in D_{k}$. It follows that if $e$ and $f$ are two idempotents in $D_{k}$ then they are isomorphic, i.e., there exists $x\in S$ such that $x^{-1}x=e$ and $xx^{-1}=f$.
Proposition F.
$\mathbb{C}D_{k}$ is an algebra.
Proof: From the definition, $\mathbb{C}D_{k}$ is a vector space. Let $s,t\in D_{k}$ be two arbitrary elements. So, the multiplication of $\lfloor s\rfloor$ and $\lfloor t\rfloor$ is given by Eq. (10). Now, it is sufficient to show that $st\in D_{k}$. As $s,t\in D_{k}$, there exist $x\in S$, such that $x^{-1}x=ss^{-1}$ and $xx^{-1}=tt^{-1}$. Now
$(st)(st)^{-1}$ $=stt^{-1}s^{-1}$
$=ss^{-1}ss^{-1}$
$=ss^{-1}$
$=x^{-1}x,$
where in the 2nd equality, we have used the condition in Eq. (10). It is clear from the above equation that $(st)\mathcal{D}t$ and hence $st\in D_{k}$. This completes the proof. $\Box$
A subset $G\subseteq S$ is called a subgroup of a semigroup $S$ if $G$ is a group with respect to the operations in $S$. A subgroup $G$ of $S$ is called a maximal subgroup if it is not properly contained in some subgroup of $S$ except itself.
Proposition.
[4] Let $S$ be a finite inverse semigroup and $e\in S$ be an idempotent. Then $e$ is the identity element of a unique maximal subgroup $G_{e}\subseteq S$ given by
$G_{e}=\{s\in S\quad|\quad ss^{-1}=s^{-1}s=e\},$
Proposition.
Let $e$ and $f$ be two isomorphic idempotents of a finite inverse semigroup $S$ and let $G_{e}$ and $G_{f}$ represents their maximal subgroups. Then $G_{e}$ is isomorphic to $G_{f}$.
From Proposition (F) and the multiplication rule given in Eq. (10), it follows that $\mathbb{C}_{0}[S]=\bigoplus_{k=1}^{n}\mathbb{C}D_{k}$. For each $D_{k}$, let us fix an idempotent $e_{k}$. Let $G_{e_{k}}$ denotes the maximal subgroup of $S$ at $e_{k}$. For each idempotent $e\in D_{k}$, let us fix an element $p_{e}\in S$ such that $\mathrm{dom}(p_{e})=e_{k}$ and $\mathrm{ran}(p_{e})=e$ with $p_{e_{k}}=e_{k}$. The following theorem states that $\mathbb{C}D_{k}$ is isomorphic to $M_{r_{k}}(\mathbb{C}G_{e_{k}})$ where $M_{r_{k}}(\mathbb{C}G_{e_{k}})$ is the algebra of $r_{k}\times r_{k}$ matrices over the group algebra $\mathbb{C}G_{e_{k}}$ with $r_{k}$ being the number of idempotents in $D_{k}$.
Theorem G.
[4] Let $S$ be a finite inverse semigroup and $\{D_{k}\}_{k=1}^{N}$ represents its $\mathcal{D}$-classes. Let $e_{k}$ be an arbitrary but fixed idempotent in $D_{k}$ and $G_{e_{k}}$ represents the maximal subgroup of at $e_{k}$. Then $\mathbb{C}_{0}[S]\cong\bigoplus_{k=1}^{N}M_{r_{k}}(\mathbb{C}G_{k})$. Specifically, The algebra $\mathbb{C}D_{k}$ is isomorphic to the algebra $M_{r_{k}}(\mathbb{C}G_{k})$, where the isomorphism is given by
$\varphi(\lfloor s\rfloor)=p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}E_{\mathrm{ran}(s),\mathrm{dom}(s)},$ (11)
and the inverse map is given by
$\varphi^{-1}(sE_{t,f})=\lfloor p_{t}sp^{-1}_{f}\rfloor.$ (12)
Proposition H.
Let $s\in D_{k}$. Then $p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\in G_{e_{k}}$.
Proof: We begin with the following,
$\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)^{-1}\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)$ $=p^{-1}_{\mathrm{dom}(s)}s^{-1}p_{\mathrm{ran}(s)}p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}$
$=p^{-1}_{\mathrm{dom}(s)}s^{-1}\mathrm{ran}(s)sp_{\mathrm{dom}(s)}$
$=p^{-1}_{\mathrm{dom}(s)}s^{-1}ss^{-1}sp_{\mathrm{dom}(s)}$
$=p^{-1}_{\mathrm{dom}(s)}s^{-1}sp_{\mathrm{dom}(s)}$
$=p^{-1}_{\mathrm{dom}(s)}\mathrm{dom}(s)p_{\mathrm{dom}(s)}$
$=p^{-1}_{\mathrm{dom}(s)}p_{\mathrm{dom}(s)}p^{-1}_{\mathrm{dom}(s)}p_{\mathrm{dom}(s)}$
$=p^{-1}_{\mathrm{dom}(s)}p_{\mathrm{dom}(s)}$
$=e_{k}.$
In a similar way, we can also prove that $\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)^{-1}=e_{k}$. This completes the proof. $\Box$
Proposition I.
For all $u\in G_{e_{k}}$, we have $\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a$ and $\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b$.
Proof:
${ran}(p_{a}\,u\,p^{-1}_{b})$ $=p_{a}\,u\,p^{-1}_{b}\,p_{b}\,u^{-1}\,p^{-1}_{a}$
$=p_{a}\,u\,e_{k}\,u^{-1}\,p^{-1}_{a}$
$=p_{a}\,u\,u^{-1}\,p^{-1}_{a}$
$=p_{a}\,e_{k}\,p^{-1}_{a}$
$=p_{a}\,p^{-1}_{a}\,p_{a}\,p^{-1}_{a}$
$=p_{a}\,p^{-1}_{a}$
$=a.$
In a similar way, we can prove $\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b$. This completes the proof. $\Box$
Steinberg’s isomorphism provides a way to construct a complete set of inequivalent irreducible representations of the contracted algebra $\mathbb{C}_{0}[S]$ by using complete sets of inequivalent irreducible representations of the maximal subgroups of the finite inverse semigroup $S$.
Theorem J.
[4] Let $S$ be a finite inverse semigroup and $\{D_{k}\}_{k=1}^{N}$ represents its $\mathcal{D}$-classes. Let $e_{k}$ be an arbitrary but fixed idempotent in $D_{k}$ and $G_{e_{k}}$ represents the maximal subgroup of at $e_{k}$. Let $\mathrm{Irr}(G_{k})$ denotes a complete set of inequivalent irreducible representations of the maximal subgroup $G_{e_{k}}$. For $\rho_{k}\in\mathrm{Irr}(G_{k})$, define a representation $\overline{\rho}_{k}$ on $M_{r_{k}}(\mathbb{C}G_{k})$ by $\overline{\rho}_{k}(gE_{a,b}):=E_{a,b}\otimes\rho_{k}(g)$ with the linear extension to the whole algebra. Now define $\overline{\rho}_{k}$ on the algebra $\bigoplus_{k=1}^{N}M_{r_{k}}(\mathbb{C}G_{k})$ by defining its action on the summands other than $k$ as zero. Then the set of $\overline{\rho}_{k}$ defined in this way forms a complete set of inequivalent irreducible representation of $\mathbb{C}_{0}[S]$.
II.2 CJKS Isomorphism and completely positive maps
Let $\mathcal{H}_{1}=\mathbb{C}^{n}$ and $\mathcal{H}_{2}=\mathbb{C}^{m}$ be two finite-dimensional Hilbert spaces and $B(\mathcal{H}_{1})$ and $B(\mathcal{H}_{2})$ respectively represent the space of bounded linear operators on them. Let $\{e_{ij}\}_{i,j=1}^{n}$ and $\{f_{ij}\}_{i,j=1}^{m}$ respectively represents the complete set of matrix units for $B(\mathcal{H}_{1})$ and $B(\mathcal{H}_{2})$. Let $\Phi:B(\mathcal{H}_{1})\rightarrow B(\mathcal{H}_{2})$ be a linear map. Then the Choi matrix [5, 6, 1, 7, 8] of $\Phi$ is defined as
$C_{\Phi}=\sum_{{i,j}=1}^{n}e_{ij}\otimes\Phi(e_{ij})\in B(\mathcal{H}_{1})\otimes B(\mathcal{H}_{2}).$ (13)
The map $\Phi\mapsto C_{\Phi}$ is known as the CJKS isomorphism with the inverse given by
$\Phi(A)=\operatorname{tr}_{1}[(A^{\tau}\otimes\mathbb{I})C_{\Phi}],$ (14)
where $\tau$ represent transposition with respect to the usual canonical basis of $\mathcal{H}_{1}$ and $\mathbb{I}$ represents the identity operator on $\mathcal{H}_{2}$. The Eq. (14) will be referred as the Choi inversion formula throughout the paper.
A map $\Phi:B(\mathcal{H}_{1})\rightarrow B(\mathcal{H}_{2})$ is called positive if it maps positive elements of $B(\mathcal{H}_{1})$ to the positive elements of $B(\mathcal{H}_{2})$. Now using $\Phi$, one can construct a map $\mathrm{id}_{k}\otimes\Phi:M_{k}(\mathbb{C})\otimes B(\mathcal{H}_{1})\rightarrow M_{k}(\mathbb{C})\otimes B(\mathcal{H}_{2})$, where $M_{k}(\mathbb{C})$ is the space of $k\times k$ matrices whose entries are complex numbers.
Definition K.
The map $\Phi:B(\mathcal{H}_{1})\rightarrow{B}(\mathcal{H}_{2})$ is called $k$-positive if $\mathrm{id}_{k}\otimes\Phi:M_{k}(\mathbb{C})\otimes B(H_{1})\rightarrow M_{k}(\mathbb{C})\otimes B(H_{2})$ is positive. If $\Phi$ is $k$-positive for all $k\in\mathbb{N}$, then $\Phi$ is called completely positive.
From the definition it is obvious that a complete positive map is a positive map, however the converse is not always true. The transposition map on matrices is an example of a map which is positive but not completely positive. The following theorem provides necessary and sufficient condition for a map to be completely positive.
Theorem L (Choi Theorem).
[5, 6, 1, 7, 8]. A map $\Phi:B(\mathcal{H}_{1})\rightarrow B(\mathcal{H}_{2})$ is completely positive if and only if the Choi matrix $C_{\Phi}=\sum_{{i,j}=1}^{n}e_{ij}\otimes\Phi(e_{ij})$ is positive semidefinite.
II.3 Space of Super-maps
We consider the Hilbert spaces $\mathcal{H}_{1}$, $\mathcal{H}_{2}$, $\mathcal{H}_{3}$ and $\mathcal{H}_{4}$ to be finite dimensional. Let $B(\mathcal{H}_{i})$ represents the space of bounded linear operators on the Hilbert space $\mathcal{H}_{i}$, for $i=1,2,3,4$. We represent by $L(B(\mathcal{H}_{i}),B(\mathcal{H}_{j}))$ the space of linear operators from $B(\mathcal{H}_{i})$ to $B(\mathcal{H}_{j})$ and by $L[L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2})),L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))]$ we represent the space of linear operators from $L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2}))$ to $L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))$. A linear map $\Theta:L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2}))\rightarrow L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))$ is called a super-map.
II.3.1 Choi-type theorem for Super-maps
[…]
$\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}(A)$ $=\operatorname{tr}({e_{ij}^{\mathcal{H}_{1}}}^{\dagger}A)e^{\mathcal{H}_{2}}_{kl},$ (15)
where $\dagger$” denotes the hermitian conjugate.
The spaces $L\left[L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2})),L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))\right]$ and $L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2}))\otimes L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))$ are isomorphic as vector spaces under the following identification
$\Theta\leftrightarrow\sum_{ijkl}\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}\otimes\Theta(\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}).$ (16)
It should be noted that the object $\Lambda_{\Theta}:=\sum_{ijkl}\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}\otimes\Theta(\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl})\in L(B(\mathcal{H}_{1}),B(\mathcal{H}_{2}))\otimes L(B(\mathcal{H}_{3}),B(\mathcal{H}_{4}))$ has the mathematical structure similar to Choi matrices (see Ref.[10])
III Main Results
Let $G$ be a finite group with cardinality $|G|$, and $\mathbb{C}[G]$ denotes the group algebra of $G$ over the complex field $\mathbb{C}$. A generic element $a\in\mathbb{C}[G]$ has the form $a=\sum_{i=1}^{|G|}\alpha_{i}g_{i}$. So, under the identification $a\leftrightarrow(\alpha_{1},\alpha_{2},....,\alpha_{|G|})$, the two spaces $\mathbb{C}[G]$ and $L(\mathbb{C}_{0}[G],\mathbb{C})$ are isomorphic as vector spaces. Now there is a natural multiplication defined on $\mathbb{C}[G]$, induced by the group multiplication, namely
$a_{1}\cdot a_{2}:=\sum_{i,j=1}^{|G|}f_{1}(g_{i})f_{2}(g_{j})g_{i}g_{j}.$ (17)
Using the fact that in the group multiplication table, along a row or column each element of the group appears once and only once, the above sum can be rewritten as
$a_{1}\cdot a_{2}$ $:=\sum_{k=1}^{|G|}\sum_{i=1}^{|G|}f_{1}(g_{i})f_{2}(g_{i}^{-1}g_{k})g_{k}$
$=\sum_{k=1}^{|G|}(f_{1}*f_{2})(g_{k})g_{k},$ (18)
where $(f_{1}*f_{2})(g_{k}):=\sum_{i=1}^{|G|}f_{1}(g_{i})f_{2}(g_{i}^{-1}g_{k})$ is the convolution of $f_{1}$ and $f_{2}$. This shows that the convolution algebra $L(\mathbb{C}_{0}[G],\mathbb{C})$ is isomorphic to the group algebra $\mathbb{C}_{0}[G]$. The above fact can be generalized as follows:
Theorem A.
[…]
See Appendix (B) for the proof. Now we generalize the above theorem in the context of contracted semigroup algebra. We begin by defining convolution in the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to an arbitrary algebra $\mathcal{A}$.
Definition 1.
We define the convolution of two maps $\mathrm{\Phi}$ and $\mathrm{\Phi^{\prime}}$ in the space $L\big{(}\mathbb{C}_{0}[S],\mathcal{A}\big{)}$ as follows:
$\mathrm{\Phi}*\mathrm{\Phi^{\prime}}(s_{k}):=\begin{cases}\sum_{i,j}\mathrm{\Phi}(s_{i})\mathrm{\Phi^{\prime}}(s_{j})&\text{if}\ s_{k}\in\mathrm{Im}(\cdot),\\ 0&\text{if}\ s_{k}\notin\mathrm{Im}(\cdot).\end{cases}$ (19)
for all $s\in S\setminus z$. Where, the sum is over the pair $(i,j)$ such that $s_{i}\cdot s_{j}=s_{k}$ and $\mathrm{Im}(\cdot)$ represents the image of the multiplication map $\cdot:S\times S\rightarrow S$ of the semigroup $S$.
Theorem 1.
Let $S$ be a semigroup with zero element $z$. We denote the contracted algebra of $S$ by $\mathbb{C}_{0}[S]$. Then the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ is isomorphic to $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ with the convolution product being defined by Eq. (19).
Proof: Let us consider a linear map $\mathrm{\Phi}:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$, where $\mathcal{A}$ is an arbitrary algebra. We denote the space of all such maps by $L\big{(}\mathbb{C}_{0}[S],\mathcal{A}\big{)}$. Again, it can be shown that the spaces $L\big{(}\mathbb{C}_{0}[S],\mathcal{A}\big{)}$ and $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ are isomorphic under the following identification:
$\mathrm{\Phi}\leftrightarrow\sum_{s_{i}\neq 0}s_{i}\otimes\mathrm{\Phi}(s_{i}).$ (20)
With the convolution product defined in Eq. (19), the space $L\big{(}\mathbb{C}_{0}[S],\mathcal{A}\big{)}$ is a convolution algebra. We have the tensor product algebra $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ where the multiplication rule comes from the individual algebras $\mathbb{C}_{0}[S]$ and $\mathcal{A}$. We now show that the identification in Eq. (20) is an algebra isomorphism. Let $x$ and $x^{\prime}$ be two arbitrary elements in $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ and let $\mathrm{\Phi}$ and $\mathrm{\Phi}^{\prime}$ be the corresponding maps in $L(\mathbb{C}_{0}[S],\mathcal{A})$. Now we have
$x\cdot x^{\prime}$ $=\bigg{(}\sum_{s_{i}\neq 0}s_{i}\otimes\mathrm{\Phi}(s_{i})\bigg{)}\cdot\bigg{(}\sum_{s_{j}\neq 0}s_{j}\otimes\mathrm{\Phi}^{\prime}(s_{j})\bigg{)}$
$=\sum_{\begin{subarray}{c}s_{i}\neq 0\\ s_{j}\neq 0\end{subarray}}s_{i}s_{j}\otimes\mathrm{\Phi}(s_{i})\mathrm{\Phi}^{\prime}(s_{j}).$ (21)
Now if we look at the term $s_{i}s_{j}$, it can either be $s_{k}\in\mathrm{Im}(\cdot)$ for some $k$ or it can be 0. In the above summation, we collect all those terms for which $s_{i}s_{j}=s_{k}\neq 0$ and rewrite Eq. (21) as follows:
$x\cdot x^{\prime}$ $=\sum_{\begin{subarray}{c}s_{k}\neq 0\\ s_{k}\in\mathrm{Im}(\cdot)\end{subarray}}s_{k}\otimes\left(\sum_{p,q}\mathrm{\Phi}(s_{p})\mathrm{\Phi}^{\prime}(s_{q})\right)+0\otimes\bigg{(}\sum_{u,v}\mathrm{\Phi}(s_{u})\mathrm{\Phi}^{\prime}(s_{v})\bigg{)},$ (22)
where the summation inside the bracket of the first term is over those pair of elements $(s_{p},s_{q})$ such that $s_{p}s_{q}=s_{k}\neq 0$ and the summation inside the bracket of the second term is over the pair $(u,v)$ such that $s_{u}\cdot s_{v}=0$. As the second term vanishes, we have
$x\cdot x^{\prime}$ $=\sum_{\begin{subarray}{c}s_{k}\neq 0\\ s_{k}\in\mathrm{Im}(\cdot)\end{subarray}}s_{k}\otimes\left(\sum_{p,q}\mathrm{\Phi}(s_{p})\mathrm{\Phi}^{\prime}(s_{q})\right)$
$=\sum_{\begin{subarray}{c}s_{k}\neq 0\\ s_{k}\in\mathrm{Im}(\cdot)\end{subarray}}s_{k}\otimes\left(\sum_{p,q}\mathrm{\Phi}(s_{p})\mathrm{\Phi}^{\prime}(s_{q})\right)+\sum_{\begin{subarray}{c}s_{k}\neq 0\\ s_{k}\notin\mathrm{Im}(\cdot)\end{subarray}}s_{k}\otimes 0$ (23)
$=\sum_{\begin{subarray}{c}s_{k}\neq 0\\ s_{k}\in S\end{subarray}}s_{k}\otimes(\mathrm{\Phi}*\mathrm{\Phi^{\prime}})(s_{k}),$ (24)
where we have used the definition of convolution (19) in the last line. This completes the proof. $\Box$
Let us now consider a $d$-dimensional Hilbert space $\mathcal{H}$. The set $S:=\{e_{ij}\}_{i,j=1}^{d}\cup\{0\}\subset\mathcal{B}(\mathcal{H})$ with “0” being the zero matrix, is a semigroup under the matrix multiplication:
$e_{ij}\cdot e_{kl}=e_{il}\delta_{jk}.$ (25)
Then the contracted algebra $\mathbb{C}_{0}[S]$ is isomorphic to the full algebra $\mathcal{B}(\mathcal{H})$ [3, 11]. Hence, we can view the matrix algebra $\mathcal{B}(\mathcal{H})$ as the contracted semigroup algebra of the semigroup $\{e_{ij}\}_{i,j=1}^{d}\cup\{0\}$. In the specific case when we consider the space $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ of linear maps from $\mathcal{B}(\mathcal{H}_{1})$ to $\mathcal{B}(\mathcal{H}_{2})$ then Theorem (1) becomes the following:
The convolution algebra $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ and the algebra $\mathcal{B}(\mathcal{H}_{1})\otimes\mathcal{B}(\mathcal{H}_{2})$ are isomorphic, with the isomorphism being the Choi-Jamiołkowski isomorphism $\Phi\mapsto\sum_{i,j}e_{ij}\otimes\Phi(e_{ij})$ and the convolution product being
$\Phi*\Phi^{\prime}(e_{ij}):=\sum_{k}\Phi(e_{ik})\Phi^{\prime}(e_{kj})$ (26)
and thereby retrieving the result of [2].
The basis $\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}$ for the space $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ is defined in Eq. (15). Now we show that $\Big{\{}\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}\Big{\}}\cup\big{\{}0\big{\}}$ forms a semigroup.
Proposition 1.
The subset $S:=\Big{\{}\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}\Big{\}}\cup\big{\{}0\big{\}}\subset L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ forms a semigroup under the convolution product defined in Eq. (26).
Proof: The convolution between two elements $\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}$ and $\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{pqrs}$ is given by
$\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}*\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{pqrs}(e^{\mathcal{H}_{1}}_{uv})$ $=\sum_{w}\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}(e^{\mathcal{H}_{1}}_{uw})\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{pqrs}(e^{\mathcal{H}_{1}}_{wv})$
$=\sum_{w}\operatorname{tr}(e^{\mathcal{H}_{1}\dagger}_{kl}e^{\mathcal{H}_{1}}_{uw})\operatorname{tr}(e^{\mathcal{H}_{1}\dagger}_{rs}e^{\mathcal{H}_{1}}_{wv})e^{\mathcal{H}_{2}}_{ij}e^{\mathcal{H}_{2}}_{pq}$
$=\sum_{w}\delta_{ku}\delta_{lw}\delta_{rw}\delta_{sv}\delta_{jp}e^{\mathcal{H}_{2}}_{iq}$
$=\delta_{ku}\delta_{sv}\delta_{lr}\delta_{jp}e^{\mathcal{H}_{2}}_{iq}$
$=\operatorname{tr}(e^{\mathcal{H}_{1}\dagger}_{ks}e^{\mathcal{H}_{1}}_{uv})e^{\mathcal{H}_{2}}_{iq}\delta_{lr}\delta_{jp}$
$=\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{iqks}(e^{\mathcal{H}_{1}}_{uv})\delta_{lr}\delta_{jp}$
which implies that $\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}*\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{pqrs}=\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{iqks}\delta_{lr}\delta_{jp}$. This completes the proof. $\Box$
Recalling that $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))$ and $L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))$ are convolution algebras with the convolution being defined as in Eq. (26), and taking in to account Proposition (1), we define convolution on $L\left[L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2})),L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))\right]$ as follows:
$\Theta_{1}\star\Theta_{2}(\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ijkl}):=\sum_{pq}\Theta_{1}(\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{ipkq})*\Theta_{2}(\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{pjql}).$ (27)
In this setting Theorem (1) becomes the following:
The convolution algebra $L\left[L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2})),L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))\right]$ with the convolution being defined by Eq. (27) and the convolution algebra $L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2}))\otimes L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))$ with the convolution defined by Eq. (26) are isomorphic under the identification $\Theta\mapsto\sum_{ijkl}\mathcal{E}^{H_{1}\rightarrow H_{2}}_{ijkl}\otimes\Theta(\mathcal{E}^{H_{1}\rightarrow H_{2}}_{ijkl})$.
A super-map $\mathrm{\Theta}:L\left(\mathcal{B}\left(\mathcal{H}_{1}\right),\mathcal{B}\left(\mathcal{H}_{2}\right)\right)\rightarrow L\left(\mathcal{B}\left(\mathcal{H}_{3}\right),\mathcal{B}\left(\mathcal{H}_{4}\right)\right)$ induces a map $T:\mathcal{B}\left(\mathcal{H}_{1}\right)\otimes\mathcal{B}\left(\mathcal{H}_{2}\right)\rightarrow\mathcal{B}\left(\mathcal{H}_{3}\right)\otimes\mathcal{B}\left(\mathcal{H}_{4}\right)$ known as “representing map” [12] at the level of Choi matrices and is given by the following,
$T(X)=C_{\mathrm{\Theta}\left(\Gamma_{X}\right)},$ (28)
where $X\in\mathcal{B}\left(\mathcal{H}_{1}\right)\otimes\mathcal{B}\left(\mathcal{H}_{2}\right)$ is the Choi matrix of $\Gamma_{X}$, and $C_{\mathrm{\Theta}\left(\Gamma_{X}\right)}$ is the Choi matrix for map $\mathrm{\Theta}\left(\Gamma_{X}\right)$. Now, under the identification: $\mathrm{\Theta}\leftrightarrow T$, we have [12, 9, 13]
$L[L\left(\mathcal{B}\left(\mathcal{H}_{1}\right),\mathcal{B}\left(\mathcal{H}_{2}\right)\right),L\left(\mathcal{B}\left(\mathcal{H}_{3}\right),\mathcal{B}\left(\mathcal{H}_{4}\right)\right)]\cong L[\mathcal{B}\left(\mathcal{H}_{1}\right)\otimes\mathcal{B}\left(\mathcal{H}_{2}\right),\mathcal{B}\left(\mathcal{H}_{3}\right)\otimes\mathcal{B}\left(\mathcal{H}_{4}\right)].$
We now prove that the map $\mathrm{\Theta}\leftrightarrow T$ preserves the respective convolutions. Let $T\leftrightarrow\Theta_{1}\star\Theta_{2}$. Then we want to show that $T=T_{1}*T_{2}$:
$T\left(e^{\mathcal{H}_{1}}_{ij}\otimes e^{\mathcal{H}_{2}}_{kl}\right)$ $=T\left(C_{\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{klij}}\right)$
$=C_{\Theta_{1}\star\Theta_{2}\left({\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{klij}}\right)}$
$=\sum_{pq}C_{\Theta_{1}\left({\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{kpiq}}\right)}C_{\Theta_{2}\left({\mathcal{E}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{2}}_{plqj}}\right)}$
$=\sum_{pq}T_{1}\Big{(}e^{\mathcal{H}_{1}}_{iq}\otimes e^{\mathcal{H}_{2}}_{kp}\Big{)}T_{2}\Big{(}e^{\mathcal{H}_{1}}_{qj}\otimes e^{\mathcal{H}_{2}}_{pl}\Big{)}$
$=T_{1}*T_{2}\Big{(}e^{\mathcal{H}_{1}}_{ij}\otimes e^{\mathcal{H}_{2}}_{kl}\Big{)}.$ (29)
With the above observation, the convolution algebras $L[L(\mathcal{B}(\mathcal{H}_{1}),\mathcal{B}(\mathcal{H}_{2})),L(\mathcal{B}(\mathcal{H}_{3}),\mathcal{B}(\mathcal{H}_{4}))]$ and $L[\mathcal{B}\left(\mathcal{H}_{1}\right)\otimes\mathcal{B}\left(\mathcal{H}_{2}\right),\mathcal{B}\left(\mathcal{H}_{3}\right)\otimes\mathcal{B}\left(\mathcal{H}_{4}\right)]$ are isomorphic.
III.1 Fourier transform of maps
Let $G$ be a finite group and $f:G\rightarrow\mathbb{C}$ be a complex-valued function on $G$. Let $\rho:G\rightarrow M_{d_{\rho}}(\mathbb{C})$ be a representation of the group $G$ of degree $d_{\rho}$. The Fourier transform of the function $f$ w.r.t the representation $\rho$ is given by:
$\widehat{f}(\rho)=\sum_{g\in G}f(g)\rho(g).$ (30)
The inverse Fourier transform is given by
$f(g)=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}\big{(}\widehat{f}(\rho)\rho(g^{-1})\big{)}.$ (31)
If instead of the complex valued function $f:G\rightarrow\mathbb{C}$, we take a map $\Phi:G\rightarrow\mathcal{A}$, where $\mathcal{A}$ is an arbitrary algebra, then it is natural to define the Fourier transform of $\Phi$ w.r.t a representation $\rho:G\rightarrow M_{d_{\rho}}(\mathbb{C})$ as follows:
$\widehat{\Phi}(\rho)=\sum_{g\in G}\rho(g)\otimes\Phi(g)\in M_{d_{\rho}}(\mathbb{C})\otimes\mathcal{A}.$ (32)
The inverse Fourier transform is given by
$\Phi(g)=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{-1})\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]},$ (33)
where $\operatorname{tr}_{d_{\rho}}$ represents partial trace over $M_{d_{\rho}}(\mathbb{C})$ and $\mathrm{Irr}(G)$ represents the complete set of inequivalent irreducible representations of $G$. The Plancherel formula is given as follows( See Appendix (A) for a proof):
$\sum_{g\in G}\Phi(g^{-1})\Psi(g)$ $=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big{[}\widehat{\Phi}(\rho)\widehat{\Psi}(\rho)\big{]}.$ (34)
Using the definition of convolution used in Theorem (A), one can prove that the Fourier transform of the convolution of two maps on a finite group is the product of their respective Fourier transform.
III.1.1 Fourier transform of maps from a contracted algebra to an arbitrary algebra
Let $S$ be a finite semigroup and $z\in S$ be the zero element. We define the Fourier transform of a linear map $\Phi:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$ w.r.t a representation $\rho:\mathbb{C}_{0}[S]\rightarrow M_{d_{\rho}}(\mathbb{C})$ as
$\widehat{\Phi}(\rho)=\sum_{\begin{subarray}{c}s\neq 0\\ s\in S\end{subarray}}\rho(s)\otimes\Phi(s)\in M_{d_{\rho}}(\mathbb{C})\otimes\mathcal{A}.$ (35)
Now we prove the convolution theorem for the Fourier transform of maps on the contracted semigroup algebras.
Proposition 2.
Let $\mathbb{C}_{0}[S]$ be the contracted semigroup algebra of a semigroup $S$ and $\mathcal{A}$ be an arbitrary algebra. Then for maps $\Phi:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$ and $\Phi^{\prime}:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$, and a representation $\rho:\mathbb{C}_{0}[S]\rightarrow M_{d_{\rho}}(\mathbb{C})$ we have
$\widehat{\Phi*\Phi^{\prime}}(\rho)=\widehat{\Phi}(\rho)\widehat{\Phi^{\prime}}(\rho).$ (36)
Proof: We begin with the product of the Fourier transforms $\widehat{\Phi}(\rho)$ and $\widehat{\Phi^{\prime}}(\rho)$, i.e.,
$\widehat{\Phi}(\rho)\widehat{\Phi^{\prime}}(\rho)$ $=\bigg{(}\sum_{s\in S,s\neq 0}\rho(s)\otimes\mathrm{\Phi}(s)\bigg{)}\cdot\bigg{(}\sum_{t\in S,t\neq 0}\rho(t)\otimes\mathrm{\Phi}^{\prime}(t)\bigg{)}$
$=(\rho\otimes\mathrm{id})\sum_{\begin{subarray}{c}s,t\in S\\ s,t\neq 0\end{subarray}}st\otimes\Phi(s)\Phi^{\prime}(t)$
$=(\rho\otimes\mathrm{id})\sum_{\begin{subarray}{c}u\neq 0\\ u\in S\end{subarray}}u\otimes(\mathrm{\Phi}*\mathrm{\Phi^{\prime}})(u)$
$=\widehat{\Phi*\Phi^{\prime}}(\rho),$
where the 2nd equality follows from the fact that $\rho$ is a representation, and the 3rd equality follows from the arguments used in Eqs. (21), (22), (23) and (24). This completes the proof. $\Box$
Now one can ask if there is a one-to-one correspondence between $\Phi$ and $\widehat{\Phi}(\rho)$. In general there is no such correspondence, however if we consider a complete set of irreducible representations $\rho$ of $\mathbb{C}_{0}[S]$, then Wedderburn-Artin theorem helps us to built a one-to-one correspondence. The inverse semigroup algebra $\mathbb{C}[S]$ is semisimple [14]. Now it can be proved that $\mathbb{C}_{0}[S]$ is semisimple if and only if $\mathbb{C}[S]$ is semisimple [3]. Wedderburn isomorphism states that $\mathbb{C}_{0}[S]$ is isomorphic to $\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C})$, with $\mathrm{Irr}(\mathbb{C}_{0}[S])$ being the complete set of inequivalent irreducible representations of the contracted algebra $\mathbb{C}_{0}[S]$. The semigroup algebra $\mathbb{C}_{0}[S]$ is isomorphic to the convolution algebra $L(\mathbb{C}_{0}[S],\mathbb{C})$ implying that $L(\mathbb{C}_{0}[S],\mathbb{C})\cong\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C})$. In particular, the isomorphism is given by $\Phi\mapsto\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}\left(\sum_{s\in S\setminus\{0\}}\Phi(s)\rho(s)\right)$. This simple observation generalizes to $L(\mathbb{C}_{0}[S],\mathcal{A})\cong\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C})\otimes\mathcal{A}$ with $\Phi\mapsto\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}\left(\sum_{s\in S\setminus\{0\}}\rho(s)\otimes\Phi(s)\right)$. Equivalently, we can say that there is one-to-one correspondence between $\Phi$ and $\bigoplus_{\rho\in\mathcal{X}}\widehat{\Phi}(\rho)$.
III.2 Fourier Inversion for finite inverse semigroup
In this subsection, we will derive the Fourier inversion formula for finite inverse semigroup. The main idea behind the proof is the following: Steinberg’s isomorphism provides a way to construct a complete set of inequivalent irreducible representations of the contracted algebra of a finite inverse semigroup in terms of the complete set or inequivalent irreducible representations of the maximal subgroups of the $D$-classes of the inverse semigroup (see Theorem (J)). Then the Fourier inversion formula for the maximal subgroups helps to provide the Fourier inversion for the inverse semigroup. The Fourier inversion formula for complex valued functions on the semigroup algebra of a finite inverse semigroup is derived in [15]. Here, following a similar technique, we derive the Fourier inversion formula for matrix-valued functions on the contracted algebra of a finite inverse semigroup.
Theorem 2.
Let $\mathbb{C}_{0}[S]$ be the contracted semigroup algebra of an inverse semigroup $S$. Let $\mathrm{Irr}(\mathbb{C}_{0}[S])$ denotes a complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$. Let $\{D_{k}\}_{k=1}^{n}$ represents the set of $\mathcal{D}$-classes of $S$, $r_{k}$ denotes the number of idempotent elements in $D_{k}$ and $G_{e_{k}}$ represents the maximal subgroup at an arbitrarily chosen idempotent element $e_{k}\in D_{k}$. Then the inversion formula for the Fourier transform $\widehat{\Phi}(\rho)=\sum_{s\in S,s\neq 0}\rho(s)\otimes\Phi(s)=\sum_{s\in S,s\neq 0}\rho(\lfloor s\rfloor)\otimes\widetilde{\Phi}(\lfloor s\rfloor)$ of a linear map $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ is given by
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\big{[}(\sigma(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]},$ (37)
where $d_{\sigma}$ is the dimension of the representation $\sigma$ and the linear map $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ is related to the map $\Phi$ by $\Phi(s)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\mu(s,t)\widetilde{\Phi}(\lfloor t\rfloor)$ and $\widetilde{\Phi}(\lfloor s\rfloor)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\Phi(t)$.
Proof: We denote the complete set of irreducible representations of $\mathbb{C}_{0}[S]$ constructed in the way mentioned in Theorem (J) by $\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$. Specifically, the representation $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$ is given by the following equation:
$\overline{\rho}_{k}(\lfloor s\rfloor)=\begin{cases}E_{\mathrm{ran}(s),\mathrm{dom}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)})&\text{when}\quad s\in D_{k},\\ 0&\text{otherwise}.\end{cases}$ (38)
Notice that $\sum_{s\in S,s\neq 0}s\otimes\Phi(s)\in\mathbb{C}_{0}[S]\otimes M_{n}(\mathbb{C})$ can also be expressed in terms of the groupoid basis as $\sum_{s\in S,s\neq 0}\lfloor s\rfloor\otimes\widetilde{\Phi}(\lfloor s\rfloor)$. The linear maps $\widetilde{\Phi}(\lfloor s\rfloor)$ and $\Phi(s)$ are related to each other by the following relations
$\Phi(s)$ $=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\mu(s,t)\widetilde{\Phi}(\lfloor t\rfloor),$ (39)
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\Phi(t).$ (40)
As the Fourier transform of $\Phi$ is $\widehat{\Phi}(\rho)=(\rho\otimes\mathrm{id})\sum_{s\in S,s\neq 0}s\otimes\Phi(s)$, we have
$\widehat{\Phi}(\rho)=\sum_{s\in S,s\neq 0}\rho(s)\otimes\Phi(s)=\sum_{s\in S,s\neq 0}\rho(\lfloor s\rfloor)\otimes\widetilde{\Phi}(\lfloor s\rfloor).$ (41)
Now for $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$,
$\widehat{\Phi}(\overline{\rho}_{k})=\sum_{s\in D_{k}}E_{\mathrm{ran}(s),\mathrm{dom}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)})\otimes\widetilde{\Phi}(\lfloor s\rfloor).$ (42)
Notice that the Fourier transform $\widehat{\Phi}(\bar{\rho}_{k})$ is a matrix indexed by the idempotents of $S$. For idempotents $a,b\in D_{k}$, the $(a,b)$-th element of $\widehat{\Phi}(\bar{\rho}_{k})$ is given by
$\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ $=\sum_{\begin{subarray}{c}s\in D_{k}\\ \mathrm{ran}(s)=a\\ \mathrm{dom}(s)=b\end{subarray}}\rho_{k}(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)})\otimes\widetilde{\Phi}(\lfloor s\rfloor).$ (43)
We notice from Proposition (H) that $u:=p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}$ is an element of $G_{e_{k}}$. We can invert the relation $u=p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}$ as follows: $u=p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\iff p_{\mathrm{ran}(s)}up^{-1}_{\mathrm{dom}(s)}=p_{\mathrm{ran}(s)}p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}p^{-1}_{\mathrm{dom}(s)}\iff p_{\mathrm{ran}(s)}up^{-1}_{\mathrm{dom}(s)}=\mathrm{ran}(s)\,s\,\mathrm{dom}(s)=ss^{-1}ss^{-1}s=s$. As $\mathrm{ran}(s)=a$ and $\mathrm{dom}(s)=b$, we have, $s=p_{a}\,u\,p^{-1}_{b}$. With this observation, we can write the summation in Eq. (43) as sum over those elements $u\in G_{e_{k}}$ such that $\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a$ and $\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b$:
$\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ $=\sum_{\begin{subarray}{c}u\in G_{e_{k}}\\ :\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a,\\ \mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b\end{subarray}}\rho_{k}(u)\otimes\widetilde{\Phi}(\lfloor p_{a}\,u\,p^{-1}_{b}\rfloor).$ (44)
Using Proposition (I), we can rewrite Eq. (44) as follows:
$\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ $=\sum_{\begin{subarray}{c}u\in G_{e_{k}}\end{subarray}}\rho_{k}(u)\otimes\widetilde{\Phi}(\lfloor p_{a}\,u\,p^{-1}_{b}\rfloor).$ (45)
We can immediately identify $\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ in Eq. (45) as the Fourier transformation of the map $\widetilde{\Phi}(\lfloor p_{a}\,(\cdot)\,p^{-1}_{b}\rfloor)$ on the maximal subgroup $G_{e_{k}}$. Hence, we use Eq. (33) for the Fourier inversion and arrive at the following:
$\widetilde{\Phi}(\lfloor s\rfloor)=\widetilde{\Phi}(\lfloor p_{a}\,u\,p^{-1}_{b}\rfloor)$ $=\frac{1}{|G_{e_{k}}|}\sum_{\rho_{k}\in\mathrm{Irr(G)}}d_{\rho_{k}}\operatorname{tr}_{d_{\rho_{k}}}\big{[}(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\big{]},$ (46)
where $a=\mathrm{ran}(s)$ and $b=\mathrm{dom}(s)$. From Eq. (42), we have
$\widehat{\Phi}(\bar{\rho}_{k})=\sum_{i,j\in D_{k}}E_{i,j}\otimes\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{i,j},$ (47)
where $i$ and $j$ are idempotents in $D_{k}$. A quick calculation shows that
$\operatorname{tr}_{d_{\rho_{k}}}\big{[}(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\big{]}$ $=(\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}})\big{[}(E_{b,a}\otimes\rho_{k}(u^{-1})\otimes\mathbb{I})\widehat{\Phi}(\bar{\rho}_{k})\big{]}.$ (48)
As $s\in D_{k}$, from Eq. (38), we have
$\overline{\rho}_{k}(\lfloor s\rfloor)$ $=E_{\mathrm{ran}(s),\mathrm{dom}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)}).$ (49)
Notice that $\mathrm{ran}(s)=\mathrm{dom}(s^{-1})$ and $\mathrm{dom}(s)=\mathrm{ran}(s^{-1})$. Recalling that $u^{-1}=p^{-1}_{\mathrm{dom}(s)}\,s^{-1}\,p_{\mathrm{ran}(s)}$, we have
$\overline{\rho}_{k}(\lfloor s^{-1}\rfloor)$ $=E_{\mathrm{ran}(s^{-1}),\mathrm{dom}(s^{-1})}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s^{-1})}\,s^{-1}\,p_{\mathrm{dom}(s^{-1})})$
$=E_{\mathrm{dom}(s),\mathrm{ran}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{dom}(s)}\,s^{-1}\,p_{\mathrm{ran}(s)})$
$=E_{b,a}\otimes\rho_{k}(u^{-1}).$ (50)
With the above observation, Eq. (48) becomes the following:
$\operatorname{tr}_{d_{\rho_{k}}}\big{[}(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\big{]}$ $=(\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}})\big{[}(\overline{\rho}_{k}(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\bar{\rho}_{k})\big{]}.$ (51)
Notice that $\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}}=\operatorname{tr}_{d_{\overline{\rho}_{k}}}$. Using Eq. (51) in Eq. (46), we have
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\frac{1}{|G_{e_{k}}|}\sum_{\rho_{k}\in\mathrm{Irr(G)}}d_{\rho_{k}}\operatorname{tr}_{d_{\overline{\rho}_{k}}}\big{[}(\overline{\rho}_{k}(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\bar{\rho}_{k})\big{]}.$ (52)
Let $\mathrm{Irr}(\mathbb{C}_{0}[S])$ be an arbitrary complete set of inequivalent irreducible representation of $\mathbb{C}_{0}[S]$. Let $\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])$. Then $\sigma$ must be equivalent to an element $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$, i.e there exists an invertible transformation $H$ such that $\overline{\rho}_{k}(s)=H^{-1}\sigma(s)H$ for all $s\in S$. From Eq. (41), we have
$\widehat{\Phi}(\rho)$ $=\sum_{s\in S,s\neq 0}\overline{\rho}_{k}(\lfloor s\rfloor)\otimes\widetilde{\Phi}(\lfloor s\rfloor)$
$=\sum_{s\in S,s\neq 0}H^{-1}\sigma(\lfloor s\rfloor)H\otimes\widetilde{\Phi}(\lfloor s\rfloor)$
$=(H^{-1}\otimes\mathbb{I})\widehat{\Phi}(\sigma)(H\otimes\mathbb{I}).$ (53)
Using the above relation in Eq. (52), we get
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\frac{1}{|G_{e_{k}}|}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\rho_{k}}\operatorname{tr}_{d_{\overline{\rho}_{k}}}\big{[}(H^{-1}\sigma(\lfloor s^{-1}\rfloor)H\otimes\mathbb{I})(H^{-1}\otimes\mathbb{I})\widehat{\Phi}(\sigma)(H\otimes\mathbb{I})\big{]}$
$=\frac{1}{|G_{e_{k}}|}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\rho_{k}}\operatorname{tr}_{d_{\sigma}}\big{[}(HH^{-1}\sigma(\lfloor s^{-1}\rfloor)HH^{-1}\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]}$
$=\frac{1}{|G_{e_{k}}|}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\rho_{k}}\operatorname{tr}_{d_{\sigma}}\big{[}(\sigma(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]}.$ (54)
We notice that by construction, the dimension of the induced representation $d_{\overline{\rho}_{k}}=r_{k}d_{\rho_{k}}$. As $\overline{\rho}_{k}$ is equivalent to the representation $\sigma$, their dimension must be the same, i.e., $d_{\sigma}=d_{\overline{\rho}_{k}}=r_{k}d_{\rho_{k}}$ implying $d_{\rho_{k}}=\frac{d_{\sigma}}{r_{k}}$. With this observation, we have the following inversion formula
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\big{[}(\sigma(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]}.$ (55)
This completes the proof. $\Box$
Now we focus on the finite dimensional matrix algebra $M_{m}(\mathbb{C})$. It can be verified that the set $S_{M}:=\{e_{ij}\}_{i,j=1}^{m}\cup\{0\}$ of matrix unit along with the zero matrix form an inverse semigroup and the matrix algebra $M_{m}(\mathbb{C})$ is the contracted semigroup algebra, i.e., $\mathbb{C}_{0}[S_{M}]=M_{m}(\mathbb{C})$. The idempotent elements of $S_{M}$ are $e_{ii}$ for all $i\in\{1,2,...,m\}$ (and of course 0). The inverse of an element $e_{ij}$ is the $e_{ji}$ i.e., $e_{ij}^{-1}=e_{ji}$. It can also be verified that the maximal subgroup at an idempotent $e_{ii}$ is $G_{e_{ii}}=\{e_{ii}\}$.
Proposition B.
The $D$-classes of the inverse semigroup $S_{M}$ are $D_{0}:=\{0\}$ and $D_{1}:=\{e_{ij}\}_{i,j=1}^{m}$.
Proof: Let $e_{ij}\mathcal{D}~0$. Then there exist $e_{pq}$ such that $e_{qp}e_{pq}=0$ and $e_{pq}e_{qp}=e_{ij}e_{ji}$. However For a matrix unit $e_{pq}$, we always have $e_{qp}e_{pq}\neq 0$. Hence the $D$-class containing the 0``0" matrix has the element 0``0" only. Now, we have the element $e_{ki}$ which satisfies $e_{ki}^{-1}e_{ki}=e_{ij}e_{ij}^{-1}$ and $e_{ki}e_{ki}^{-1}=e_{kl}e_{kl}^{-1}$. Hence we have $e_{ij}\mathcal{D}e_{kl}$ for arbitrary $i,j,k,l\in\{1,2,...,m\}$. This completes the proof. $\Box$
Corollary 1.
Let us consider the matrix unit inverse semigroup $S_{M}:=\{e_{ij}\}_{i,j=1}^{m}\cup\{0\}$. Then the Fourier inversion formula in Eq. (37) becomes the Choi inversion formula in Eq. (14).
Proof: Let us start by recalling that two elements $s$ and $t$ belonging to an inverse semigroup $S$ satisfies $s\leq t$ if and only if there exists an idempotent $e\in S$ such that $s=et$ (see Definition (D)). Trivially we have $0\leq e_{ij}$ for all $i,j\in\{1,2,...,m\}$. Now, let $e_{ij}\leq e_{kl}$. Then there exists an idempotent $e_{pp}$ such that $e_{ij}=e_{pp}e_{kl}=e_{pl}\delta_{pk}$. As $e_{ij}\neq 0$, we must have $p=k$ which implies that $e_{ij}=e_{kl}$, which further implies that the only element $x\in\{e_{ij}\}_{i,j=1}^{m}$ that satisfies $e_{ij}\leq x$ is $x=e_{ij}$. From Eq. (8), we have that the groupoid basis corresponding to $e_{ij}$ is $e_{ij}$ itself, i.e., $\lfloor e_{ij}\rfloor=e_{ij}$ and from Eq. (40), we have $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$. As $D_{1}$ contains $m$ idempotents, we have $r_{1}=m$ and $G_{e_{ii}}$ being a group with only one element in it, we have $|G_{e_{ii}}|=1$. Using all these information in Eq. (37), we have
$\widetilde{\Phi}(\lfloor e_{ij}\rfloor)$ $=\Phi(e_{ij})=\frac{1}{n}\sum_{\sigma\in\mathrm{Irr}(M_{n}(\mathbb{C}))}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\big{[}(\sigma(e_{ji})\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]}.$ (56)
Now, up to equivalence, there is exactly one irreducible representation of the full matrix algebra $M_{m}(\mathbb{C})$ [16]. We can “take” this representation to be the identity map $\mathrm{id}$" on the matrix algebra $M_{m}(\mathbb{C})$ (under the identification of $M_{m}(\mathbb{C})$ with $\mathrm{End}(\mathbb{C}^{m})$). Now using the fact that the dimension of this representation is $m$, i.e., $d_{\mathrm{id}}=m$ and the Fourier transform of the map $\Phi$ with respect to the representation $\mathrm{id}$" is the Choi matrix of $\Phi$, i.e., $\widehat{\Phi}(\mathrm{id})=C_{\Phi}$, we have
$\Phi(e_{ij})=\operatorname{tr}_{m}\big{[}(e_{ji}\otimes\mathbb{I})C_{\Phi}\big{]},$ (57)
or equivalently for an arbitrary $x\in M_{n}(\mathbb{C})$, we have
$\Phi(x)=\operatorname{tr}_{m}\big{[}(x^{\tau}\otimes\mathbb{I})C_{\Phi}\big{]},$ (58)
where $\tau$ represents the transposition map. This completes the proof. $\Box$
Theorem 3.
The Plancherel Formula is given by
$\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}r_{k}|G_{e_{k}}|\widetilde{\Phi}(\lfloor s^{-1}\rfloor)\widetilde{\Psi}(\lfloor s\rfloor)$ $=\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}[\widehat{\Phi}(\sigma)\widehat{\Psi}(\sigma)],$ (59)
where $r_{k}$ and $G_{e_{k}}$ are defined in Theorem (2).
Proof: Form Eq. (37), we have
$\widetilde{\Phi}(\lfloor s^{-1}\rfloor)$ $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\big{[}(\sigma(\lfloor s\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]}.$
Notice that $r_{k}$ and $|G_{e_{k}}|$ does not change as $s\in D_{k}$ if and only if $s^{-1}\in D_{k}$. Now multiplying both the sides by $\widetilde{\Psi}(\lfloor s\rfloor)$ and summing over $s$, we have
$\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}r_{k}|G_{e_{k}}|\widetilde{\Phi}(\lfloor s^{-1}\rfloor)\widetilde{\Psi}(\lfloor s\rfloor)$ $=\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\big{[}(\sigma(\lfloor s\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]}\widetilde{\Psi}(\lfloor s\rfloor)$
$=\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\big{[}\widehat{\Phi}(\sigma)(\sigma(\lfloor s\rfloor)\otimes\mathbb{I})\big{]}\widetilde{\Psi}(\lfloor s\rfloor)$
$=\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\big{[}\widehat{\Phi}(\sigma)(\sigma(\lfloor s\rfloor)\otimes\widetilde{\Psi}(\lfloor s\rfloor))\big{]}$
$=\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\Big{[}\widehat{\Phi}(\sigma)\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}(\sigma(\lfloor s\rfloor)\otimes\widetilde{\Psi}(\lfloor s\rfloor))\Big{]}$
$=\sum_{\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{d_{\sigma}}\Big{[}\widehat{\Phi}(\sigma)\widehat{\Psi}(\sigma)\Big{]}.$
This completes the proof. $\Box$
For matrix unit semigroup $S_{M}$, we have $r_{k}=m$, $|G_{e_{k}}|=1$, $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$ and $\widetilde{\Psi}(\lfloor e_{ij}\rfloor)=\Psi(e_{ij})$. Hence, the above formula reduces to
$\sum_{i,j=1}^{m}\Phi(e_{ji})\Psi(e_{ij})$ $=\operatorname{tr}_{m}\Big{[}\widehat{\Phi}(\mathrm{id})\widehat{\Psi}(\mathrm{id})\Big{]}=\operatorname{tr}_{m}(C_{\Phi}C_{\Psi}).$ (60)
III.3 Positive definite maps and Bochner’s theorem for finite inverse semigroup
Let $G$ be a group, and $\mathcal{H}=\mathbb{C}^{n}$ be a Hilbert space. A map $\Phi:G\to\mathcal{B}(\mathcal{H})$ is called positive definite if the matrix $\Phi(g^{-1}g^{\prime})]\in\mathcal{B}(\mathcal{H}^{\bigoplus|G|})$ is positive semidefinite, i.e.,
$\sum_{g,g^{\prime}\in G}\left<\Phi(g^{-1}g^{\prime})h_{g^{\prime}},h_{g}\right>\geq 0.$ (61)
for all subset $h:=\{h_{s}\}_{s\in G}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $G$. The positive definite-ness of a maps on a group is related to the positive semidefinite-ness of its Fourier transform by Bochner’s theorem. Let $G$ be a finite group and $f:G\to\mathbb{C}$ be a complex valued function. The Bochner’s theorem states that $f$ is positive definite if and only if its Fourier transform $\widehat{f}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU(G)}$, where $\mathrm{IrrU(G)}$ is the complete set of inequivalent irreducible unitary representations of the group $G$. The proof can be found in[17]. For matrix valued map $\Phi:G\to M_{n}(\mathbb{C})$, Bochner’s theorem becomes the following.
Bochner’s theorem for matrix valued maps: Let $G$ be a finite group. Then a map $\Phi:G\to M_{n}(\mathbb{C})$ is positive definite if and only if $\widehat{\Phi}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU}(G)$, where $\mathrm{IrrU}(G)$ represents a complete set of inequivalent irreducible unitary representations of $G$.
We provide a proof in Appendix (D).
Now we are in a position to formulate and prove Bochner’s theorem for finite inverse semigroup. We begin by defining positive definite maps on the contracted algebra of a finite inverse semigroup.
Definition 2.
Let $S$ be a finite inverse semigroup with $z$ being the zero element. A linear map $\Phi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$ where $\mathcal{H}=\mathbb{C}^{n}$, is called positive definite if the matrix $\Phi(s^{-1}s^{\prime})]\in\mathcal{B}(\mathcal{H}^{\bigoplus(|S|-1)})$ is positive semidefinite, i.e.,
$\sum_{\begin{subarray}{c}s,s^{\prime}\in S\\ s\neq 0,s^{\prime}\neq 0\end{subarray}}\left<\Phi(s^{-1}s^{\prime})h_{s^{\prime}},h_{s}\right>\geq 0$ (62)
for all subset $h:=\{h_{s}\}_{s\in S\setminus\{0\}}\subset\mathbb{C}^{n}$ whose elements are indexed by the nonzero elements of $S$.
The following propositions will be useful to prove our main results.
Proposition 3.
A linear map $\Phi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$ is positive definite if and only if the matrix $\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)]\in\mathcal{B}(\mathcal{H}^{\bigoplus(|S|-1)})$ is positive semidefinite, i.e.,
$\sum_{\begin{subarray}{c}s,s^{\prime}\in S\\ s\neq 0,s^{\prime}\neq 0\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)h_{s^{\prime}},h_{s}\right>\geq 0$ (63)
for all subset $h:=\{h_{s}\}_{s\in S\setminus\{0\}}\subset\mathbb{C}^{n}$ whose elements are indexed by the nonzero elements of $S$.
Proof: We start by recalling the definition of the groupoid basis $\lfloor s\rfloor$ of $\mathbb{C}_{0}[S]$ (see Eq. (8)).
$\lfloor s\rfloor=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t,s)t\in\mathbb{C}_{0}[S],$ (64)
where $t\neq 0$. The above equation implies that
$\lfloor s^{-1}\rfloor$ $=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t^{-1},s^{-1})t^{-1}=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t^{-1}t,s^{-1}s)t^{-1}=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t,s)t^{-1},$ (65)
where in the last two equalities we have used the fact that $\mu(t,s)=\mu(tt^{-1},ss^{-1})=\mu(t^{-1}t,s^{-1}s)$ (See [18]). Now, with $s,s^{\prime}\neq 0$, we have
$\sum_{\begin{subarray}{c}s,s^{\prime}\in S\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)h_{s^{\prime}},h_{s}\right$> $=\sum_{\begin{subarray}{c}s,s^{\prime}\in S\end{subarray}}\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\sum_{\begin{subarray}{c}t^{\prime}\in S\\ t^{\prime}\leq s^{\prime}\end{subarray}}\mu(t^{\prime},s^{\prime})\mu(t,s)\left<\Phi(t^{-1}t^{\prime})h_{s^{\prime}},h_{s}\right$>
$=\sum_{\begin{subarray}{c}s,s^{\prime},t,t^{\prime}\in S\\ t\leq s\\ t^{\prime}\leq s^{\prime}\end{subarray}}\mu(t^{\prime},s^{\prime})\mu(t,s)\left<\Phi(t^{-1}t^{\prime})h_{s^{\prime}},h_{s}\right$> (66)
$=\sum_{\begin{subarray}{c}t,t^{\prime}\in S\end{subarray}}\left<\Phi(t^{-1}t^{\prime})\sum_{\begin{subarray}{c}s^{\prime}\in S\\ t^{\prime}\leq s^{\prime}\end{subarray}}\mu(t^{\prime},s^{\prime})h_{s^{\prime}},\sum_{\begin{subarray}{c}s\in S\\ t\leq s\end{subarray}}\mu(t,s)h_{s}\right$>. (67)
Defining the set $h^{0}:=\{h^{0}_{t}\}$ consisting of elements $h^{0}_{t}:=\sum_{\begin{subarray}{c}s\in S\\ t\leq s\end{subarray}}\mu(t,s)h(s)$ (which can be inverted as $h_{s}=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}h^{0}_{t}$), we can rewrite the above equation as
$\sum_{\begin{subarray}{c}s,s^{\prime}\in S\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)h_{s^{\prime}},h_{s}\right$> $=\sum_{\begin{subarray}{c}t,t^{\prime}\in S\end{subarray}}\left<\Phi(t^{-1}t^{\prime})h^{0}_{t^{\prime}},h^{0}_{t}\right$>. (68)
This completes the proof. $\Box$
Proposition 4.
A linear map $\Phi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$ is positive definite if and only if the matrix $\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)]\in\mathcal{B}(\mathcal{H}^{\bigoplus|D_{k}|})$ is positive semidefinite for all $\mathcal{D}$-classes $D_{k}$, i.e.,
$\sum_{\begin{subarray}{c}s,s^{\prime}\in D_{k}\\ s\neq 0,s^{\prime}\neq 0\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)h_{s^{\prime}},h_{s}\right>\geq 0$ (69)
for all $D_{k}$ and all subset $h:=\{h_{s}\}_{s\in D_{k}}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $D_{k}$.
Proof: The multiplication rule for the groupoid basis as given by Eq. (10) makes the matrix $\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)]$ block diagonal because $\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor=0$ when $ss^{-1}\neq s^{\prime}s^{\prime-1}$. Now, if $ss^{-1}=s^{\prime}s^{\prime-1}$ then $s$ and $s^{\prime}$ belong to the same $\mathcal{D}$-class which implies that if $s$ and $s^{\prime}$ belong to different $\mathcal{D}$-classes then $ss^{-1}\neq s^{\prime}s^{\prime-1}$. Hence $\Phi$ is positive definite if and only if for all $\mathcal{D}$-classes $D_{k}$ the following relation holds:
$\sum_{\begin{subarray}{c}s,s^{\prime}\in D_{k}\\ s\neq 0,s^{\prime}\neq 0\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)h_{s^{\prime}},h_{s}\right>\geq 0$ (70)
for all subset $h:=\{h_{s}\}_{s\in D_{k}}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $D_{k}$. This completes the proof. $\Box$
Proposition 5.
(Schur Orthogonality relation for inverse semigroup:) Let $S$ be a finite inverse semigroup and $\{D_{k}\}_{k=1}^{n}$ represents its $\mathcal{D}$-classes. Let $s\in D_{k}$ and $\overline{\rho}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, where $\mathrm{IrrU}(\mathbb{C}_{0}[S])$ represents a complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$ induced by the complete set of inequivalent irreducible unitary representations of the maximal subgroups of $S$ as described in Theorem (J) . Let $a,b,c,d\in D_{k}$ be idempotents. Then we have
$\sum_{s\in D_{k}}((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}\overline{((\overline{\rho^{\prime}}_{k}(\lfloor s\rfloor))_{c,d})_{k,l}}$ $=\begin{cases}\frac{r_{k}|G_{e_{k}}|}{d_{\overline{\rho}_{k}}}\delta_{a,c}\delta_{b,d}\delta_{i,k}\delta_{j,l}&\text{when}~\overline{\rho}_{k}=\overline{\rho^{\prime}}_{k},\\ 0&\text{when}~\overline{\rho}_{k}\neq\overline{\rho^{\prime}}_{k},\end{cases}$ (71)
where $\overline{((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}}$ represents the complex conjugate of the $(i,j)$-th matrix element of the $(a,b)$-th block of $\overline{\rho}_{k}(\lfloor s\rfloor)$.
Proof: The action of the induced representation $\overline{\rho}_{k}$ on a groupoid basis is given by
$\overline{\rho}_{k}(\lfloor s\rfloor)=E_{\mathrm{ran}(s),\mathrm{dom}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)})\in M_{r_{k}}(M_{n}(\mathbb{C})).$ (72)
Notice that $u:=p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)}\in G_{e_{k}}$. Now from Eq. (72), it is clear that $\overline{((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}}=\overline{(\rho_{k}(u))_{i,j}}\delta_{a,\mathrm{ran}(s)}\delta_{b,\mathrm{dom}(s)}$. Now,
$\sum_{s\in D_{k}}((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}\overline{((\overline{\rho^{\prime}}_{k}(\lfloor s\rfloor))_{c,d})_{k,l}}$ $=\sum_{s\in D_{k}}(\rho_{k}(u))_{i,j}\delta_{a,\mathrm{ran}(s)}\delta_{b,\mathrm{dom}(s)}\overline{(\rho^{\prime}_{k}(u))_{k,l}}\delta_{c,\mathrm{ran}(s)}\delta_{d,\mathrm{dom}(s)}.$ (73)
Since the map $\vartheta:D_{k}\to M_{r_{K}}(G_{e_{k}})$ given by $s\mapsto(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)})E_{\mathrm{ran}(s),\mathrm{dom}(s)}$ is one-to-one correspondence with the inverse given by $gE_{e,f}\mapsto p_{e}gp_{f}^{-1}$, any $s\in D_{k}$ can be written as $s=p_{e}gp_{f}^{-1}$ for some pair of idempotents $(e,f)\in D_{k}$ and $g\in G_{e_{k}}$. It should be noted that $\mathrm{ran}(p_{e}gp_{f}^{-1})=e$ and $\mathrm{dom}(p_{e}gp_{f}^{-1})=f$ for all $g\in G_{e_{k}}$. A necessary condition for the terms in the summation of Eq. (73) to be non-vanishing is $a=\mathrm{ran}(s)$ and $b=\mathrm{dom}(s)$. So the possible non-vanishing contributions come from summing over those $s\in D_{k}$ which has the form $s=p_{a}gp_{b}^{-1}$. With this observation, we have
$\sum_{s\in D_{k}}((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}\overline{((\overline{\rho^{\prime}}_{k}(\lfloor s\rfloor))_{c,d})_{k,l}}$ $=\sum_{g\in G_{e_{k}}}(\rho_{k}(g))_{i,j}\overline{(\rho^{\prime}_{k}(g))_{k,l}}\delta_{c,a}\delta_{d,b}.$ (74)
Now by applying the Schur orthogonality relation for the inequivalent irreducible unitary representations $\rho_{k}$ and $\rho^{\prime}_{k}$ in Eq. (74), and using the fact that $d_{\bar{\rho}_{k}}=r_{k}d_{\rho_{k}}$, where $d_{\bar{\rho}_{k}}$ represents the dimension of the induced representation $\overline{\rho}_{k}$, we have
$\sum_{s\in D_{k}}((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}\overline{((\overline{\rho^{\prime}}_{k}(\lfloor s\rfloor))_{c,d})_{k,l}}$ $=\begin{cases}\frac{r_{k}|G_{e_{k}}|}{d_{\overline{\rho}_{k}}}\delta_{a,c}\delta_{b,d}\delta_{i,k}\delta_{j,l}&\text{when}~\overline{\rho}_{k}=\overline{\rho^{\prime}}_{k},\\ 0&\text{when}~\overline{\rho}_{k}\neq\overline{\rho^{\prime}}_{k}.\end{cases}$ (75)
This completes the proof. $\Box$
Before proceeding further, we recall from Eq. (38) that $\overline{\rho}_{k}(\lfloor s\rfloor)=E_{\mathrm{ran}(s),\mathrm{dom}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)})$ and $\overline{\rho}_{k}(\lfloor s\rfloor)=0$ if $s\notin D_{k}$. So, we have $(\overline{\rho}_{k}(\lfloor s\rfloor))^{\dagger}=E_{\mathrm{dom}(s),\mathrm{ran}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{dom}(s)}\,s^{-1}\,p_{\mathrm{ran}(s)})$ and using Eq. (50), we have $(\overline{\rho}_{k}(\lfloor s\rfloor))^{\dagger}=\overline{\rho}_{k}(\lfloor s^{-1}\rfloor)$. This will be useful in proving Bochner’s theorem for finite inverse semigroup.
Theorem 4.
(Bochner’s theorem for finite inverse semigroups:) Let $S$ be a finite inverse semigroup and $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ be a linear map. Then the linear map $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ defined by $\widetilde{\Phi}(\lfloor s\rfloor)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\Phi(t)$ is positive definite if and only if $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, where $\mathrm{IrrU}(\mathbb{C}_{0}[S])$ represents a complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$ induced by the complete set of inequivalent irreducible unitary representations of the maximal subgroups of $S$.
Proof: We begin by recalling the inversion formula for finite inverse semigroup
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{1}\big{[}(\sigma(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]},$ (76)
where $s\in D_{k}$. Since $\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, we have $\sigma(\lfloor s^{-1}\rfloor)=(\sigma(\lfloor s\rfloor))^{\dagger}$. So the above equation can be written as
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{1}\big{[}((\sigma(\lfloor s\rfloor))^{\dagger}\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]}.$ (77)
Let $s,t\in D_{k}$ are such that $s^{-1}s=tt^{-1}$. For such a pairs of $s$ and $t$, Eq. (78) becomes
$\widetilde{\Phi}(\lfloor s\rfloor\lfloor t\rfloor)$ $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{1}\big{[}(\sigma(\lfloor t\rfloor)^{\dagger}\sigma(\lfloor s\rfloor)^{\dagger}\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]}.$ (78)
Notice that the above equation remains valid even when $s^{-1}s\neq tt^{-1}$. Because in this case $\lfloor s\rfloor\lfloor t\rfloor=0$, and $\widetilde{\Phi}$ and $\sigma$ are being linear, we have $\widetilde{\Phi}(\lfloor s\rfloor\lfloor t\rfloor)=\sigma(\lfloor s\rfloor\lfloor t\rfloor)=0$. Now, for vectors $h_{s}\in\mathbb{C}^{n}$, indexed by $s\in D_{k}$, we have
$\sum_{s,t\in D_{k}}\left<\widetilde{\Phi}(\lfloor s^{-1}\rfloor\lfloor t\rfloor)h_{t},h_{s}\right$> $=\sum_{s,t\in D_{k}}\operatorname{tr}\left(\widetilde{\Phi}(\lfloor s^{-1}\rfloor\lfloor t\rfloor)h_{t}h_{s}^{\dagger}\right)$
$=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{s,t\in D_{k}}\sum_{\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}\big{[}(\sigma(\lfloor t\rfloor)^{\dagger}\sigma(\lfloor s\rfloor)\otimes h_{t}h_{s}^{\dagger})\widehat{\Phi}(\sigma)\big{]}$
$=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}\big{[}A_{k}(\sigma,h)^{\dagger}A_{k}(\sigma,h)\widehat{\Phi}(\sigma)\big{]},$ (79)
where $A_{k}(\sigma,h):=\sum_{s\in D_{k}}\sigma(\lfloor s\rfloor)\otimes h_{s}^{\dagger}$ and $h:=\{h_{s}\}_{s\in D_{k}}$. Note that Eq. (79) is true for all the $\mathcal{D}$-classes $D_{k}$. Now, let us assume that $\widehat{\Phi}(\sigma)$ is positive semidefinite for all $\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$. Since $A_{k}(\sigma,h)^{\dagger}A_{k}(\sigma,h)$ is positive semidefinite for any $D_{k}$, $\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$ and any choice of $h\subseteq\mathbb{C}^{n}$, the right hand side of Eq. (79) is non-negative. Hence, $\sum_{s,t\in D_{k}}\left<\widetilde{\Phi}(\lfloor s^{-1}\rfloor\lfloor t\rfloor)h_{t},h_{s}\right>\geq 0$ for all $\mathcal{D}$-classes $D_{k}$ implying that the map $\widetilde{\Phi}$ is positive definite. This completes the proof in one direction.
For the other direction, let us fix $\overline{\rho}_{k}^{0}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, where $\rho^{0}_{k}\in\mathrm{IrrU}(G_{e_{k}})$ and choose $h^{0}$ as the subset consisting of the vectors $h^{0}_{s}$ of the form
$h^{0}_{s}=\operatorname{tr}_{d_{\overline{\rho}_{k}^{0}}}\left(\overline{\rho}^{0}_{k}(\lfloor s\rfloor)\otimes\mathbb{I})B^{\dagger}\right),$ (80)
where $B\in M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})\otimes\mathbb{C}^{n}$ and $\operatorname{tr}_{d_{\overline{\rho}^{0}_{k}}}$ represents the trace over the space $M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})$. Writing $B$ as $B=\sum_{p}X_{p}\otimes x_{p}$, where $X_{p}\in M_{d_{\overline{\rho}^{0}_{k}}}(\mathbb{C})$ and $x_{p}\in\mathbb{C}^{n}$, we have
$(h_{s}^{0})^{\dagger}=\operatorname{tr}_{d_{\overline{\rho}^{0}_{k}}}\left((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right)^{\dagger}\otimes\mathbb{I})B\right)$ $=\sum_{c,d,p,k,l}\overline{((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right))_{dc})_{lk}}(((X_{p}))_{dc})_{lk}x_{p},$ (81)
where $\overline{((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right))_{dc})_{lk}}$ represents the complex conjugate of the $(l,k)$-th matrix element of the $(d,c)$-th block of the block matrix $\overline{\rho}^{0}_{k}(\lfloor s\rfloor)$. Now for $\sigma=\overline{\rho}_{k}$,
$A_{k}(\overline{\rho}_{k},h^{0})$ $=\sum_{s\in D_{k}}\overline{\rho}_{k}(\lfloor s\rfloor)\otimes(h^{0}_{s})^{\dagger}$
$=\sum_{a,b,i,j}\sum_{s\in D_{k}}((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}E_{ab}\otimes e_{ij}\otimes\sum_{c,d,p,k,l}\overline{((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right))_{dc})_{lk}}(((X_{p}))_{dc})_{lk}x_{p}$
$=\sum_{a,b,i,j}\sum_{c,d,p,k,l}\Bigg{(}\sum_{s\in D_{k}}\overline{((\overline{\rho}^{0}_{k}\left(\lfloor s\rfloor\right))_{dc})_{lk}}((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}\Bigg{)}((X_{p})_{dc})_{lk}E_{ab}\otimes e_{ij}\otimes x_{p}$
$=\begin{cases}\frac{r_{k}|G_{e_{k}}|}{d_{\overline{\rho}_{k}^{0}}}B&\text{when}~\overline{\rho}_{k}=\overline{\rho}_{k}^{0},\\ 0&\text{when}~\overline{\rho}_{k}\neq\overline{\rho}_{k}^{0},\end{cases}$ (82)
where in the last equality, we have used Schur orthogonality relation (Eq. (71)). We put the above form of $A_{k}(\overline{\rho}_{k},h^{0})$ in Eq. (79) to arrive at
$\sum_{s,t\in D_{k}}\left<\widetilde{\Phi}(\lfloor s^{-1}\rfloor\lfloor t\rfloor)h^{0}_{t},h^{0}_{s}\right$> $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}\big{[}A_{k}(\sigma,h^{0})^{\dagger}A_{k}(\sigma,h^{0})\widehat{\Phi}(\sigma)\big{]}$
$=\frac{1}{r_{k}|G_{e_{k}}|}d_{\overline{\rho}_{k}^{0}}\operatorname{tr}\big{[}A_{k}(\overline{\rho}_{k}^{0},h^{0})^{\dagger}A_{k}(\overline{\rho}_{k}^{0},h^{0})\widehat{\Phi}(\overline{\rho}_{k}^{0})\big{]}$
$=\frac{d_{\overline{\rho}_{k}^{0}}}{r_{k}|G_{e_{k}}|}\left(\frac{r_{k}|G_{e_{k}}|}{d_{\overline{\rho}_{k}^{0}}}\right)^{2}\operatorname{tr}\big{[}B^{\dagger}B\widehat{\Phi}(\overline{\rho}_{k}^{0})\big{]}$
$=\frac{r_{k}|G_{e_{k}}|}{d_{\overline{\rho}_{k}^{0}}}\operatorname{tr}\big{[}B^{\dagger}B\widehat{\Phi}(\overline{\rho}_{k}^{0})\big{]}.$ (83)
Now if we assume that $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ is positive definite then the left hand side of the above equation is non-negative. Since, $B$ and $\overline{\rho}_{k}^{0}$ are chosen arbitrarily, $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$. This completes the proof. $\Box$
The Stinespring dilation theorem for completely positive maps is a fundamental theorem in $C^{*}$ algebra. Stinespring dilation theorem also exists for positive definite maps from a group to a von Neumann algebra [19, 20]. Here we prove a Stinespring dilation theorem for positive definite maps on the contracted algebra $\mathbb{C}_{0}[S]$ of a finite inverse semigroup $S$.
Theorem 5.
(Stinespring Dilation) Let $S$ be a finite inverse semigroup and $\mathbb{C}_{0}[S]$ the the contracted algebra of $S$. Let $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ be a linear map. Then $\Phi$ is positive definite if and only if there exists a Hilbert Space $\mathcal{H}$, a bounded operator $V:\mathbb{C}^{n}\to\mathcal{H}$ and a *-homomorphism $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$ such that for all $s\neq 0$
$\Phi(\lfloor s\rfloor)=V^{*}\pi(\lfloor s\rfloor)V,$ (84)
with
$V^{*}V=\Phi(\mathds{1}),$ (85)
where $\mathds{1}:=\sum_{e\in E(S)}\lfloor e\rfloor$ with $E(S)$ denoting the set of non-zero idempotents of $S$, is the identity of the contracted algebra $\mathbb{C}_{0}[S]$.
Proof: The proof is essentially based on GNS construction. We consider the space $\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}$. Any generic element in $\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}$ can be written as $f=\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}\lfloor s\rfloor\otimes f_{s}$, where $f_{s}\in\mathbb{C}^{n}$. Assuming that the map $\Phi$ is positive definite, we can define a positive semidefinite sesquilinear form on $\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}$ as follows:
$\left<f,h\right>:=\sum_{\begin{subarray}{c}s,t\in S\\ s,t\neq 0\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor t\rfloor)f_{t},h_{s}\right$>. (86)
[…]
$\left<[f],[h]\right>:=\sum_{\begin{subarray}{c}s,t\in S\\ s,t\neq 0\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor t\rfloor)f_{t},h_{s}\right$>. (87)
It can be shown that the above defined inner product is well defined. With this inner product defined, the quotient space becomes an inner product space and as we are working on finite dimensions, it becomes a Hilbert space which we will be denoting by $\mathcal{H}:=\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}/\mathcal{N}$. Now define a linear map on $\mathcal{H}$ as
$\pi(\lfloor s\rfloor)[f]:=[\lfloor s\rfloor\cdot f],$ (88)
where the element $\lfloor s\rfloor\cdot f\in\mathbb{C}_{0}[S]\otimes\mathbb{C}^{n}$ is defined by $\lfloor s\rfloor\cdot f:=\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\lfloor s\rfloor\lfloor t\rfloor\otimes f_{t}$. The map $\pi$ extends linearly to a map $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$. One can simply verify that $\pi(\lfloor s\rfloor\lfloor t\rfloor)=\pi(\lfloor s\rfloor)\pi(\lfloor t\rfloor)$. Notice that due to the linearity of $\pi$, the relation $\pi(\lfloor s\rfloor\lfloor t\rfloor)=\pi(\lfloor s\rfloor)\pi(\lfloor t\rfloor)$ is consistent with the fact that $\lfloor s\rfloor\lfloor t\rfloor=0$ if $\mathrm{dom(s)}\neq\mathrm{ran}(t)$. Define another map $V:\mathbb{C}^{n}\to\mathcal{H}$ by
$V(x):=\left[\mathds{1}\otimes x\right],$ (89)
where $\mathds{1}:=\sum_{e\in E(S)}\lfloor e\rfloor$ with $E(S)$ denoting the set of non-zero idempotents of $S$, is the identity of the contracted algebra $\mathbb{C}_{0}[S]$ (see [4]). Let $V^{*}:\mathcal{H}\to\mathbb{C}^{n}$ denotes the adjoint of $V$. Then we have
$\left<V^{*}[f],x\right$> $=\left<[f],\left[\mathds{1}\otimes x\right]\right$>
$=\left<[f],\left[\sum_{e\in E(S)}\lfloor e\rfloor\otimes x\right]\right$>
$=\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\sum_{e\in E(S)}\left<\Phi(\lfloor e^{-1}\rfloor\lfloor t\rfloor)f_{t},x\right$>
$=\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\left<\Phi\left(\sum_{e\in E(S)}\lfloor e\rfloor\lfloor t\rfloor\right)f_{t},x\right$>
$=\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\left<\Phi\left(\lfloor t\rfloor\right)f_{t},x\right$>. (90)
So, the adjoint map $V^{*}$ is given by
$V^{*}[f]$ $=\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\Phi\left(\lfloor t\rfloor\right)f_{t}.$ (91)
Now using the definition of $V$ and the action of $V^{*}$ as provided above, we have:
$V^{*}\pi(\lfloor s\rfloor)V(x)=V^{*}\pi(\lfloor s\rfloor)\left[\mathds{1}\otimes x\right]=V^{*}\left[\lfloor s\rfloor\otimes x\right]=\Phi(\lfloor s\rfloor)x.$ (92)
Also we have
$V^{*}V(x)$ $=V^{*}\left[\mathds{1}\otimes x\right]=\Phi\left(\sum_{e\in E(S)}\lfloor e\rfloor\right)x=\Phi\left(\mathds{1}\right)x.$ (93)
Now we prove that $\pi$ is a *-homomorphism:
$\left<\pi(\lfloor u\rfloor)^{\dagger}[f],[h]\right$> $=\left<[f],\pi(\lfloor u\rfloor)[h]\right$>
$=\left<[f],[\lfloor u\rfloor\cdot h]\right$>
$=\left<\left[\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\lfloor t\rfloor\otimes f_{t}\right],\left[\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}\lfloor u\rfloor\lfloor s\rfloor\otimes h_{s}\right]\right$>
$=\sum_{\begin{subarray}{c}s,t\in S\\ s,t\neq 0\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor u^{-1}\rfloor\lfloor t\rfloor)f_{t},h_{s}\right$>
$=\left<\left[\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\lfloor u^{-1}\rfloor\lfloor t\rfloor\otimes f_{t}\right],\left[\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}\lfloor s\rfloor\otimes h_{s}\right]\right$>
$=\left<\pi(\lfloor u^{-1}\rfloor)[f],[h]\right$>. (94)
Hence $\pi(\lfloor u\rfloor)^{\dagger}=\pi(\lfloor u^{-1}\rfloor)$ for all $u\in S$, implying that $\pi$ is a *-homomorphism.
Now, we assume that a map linear $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ can be written as $\Phi(\lfloor s\rfloor)=V^{*}\pi(\lfloor s\rfloor)V$ for some Hilbert Space $\mathcal{H}$, a bounded operator $V:\mathbb{C}^{n}\to\mathcal{H}$ and a *-homomorphism $\pi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$. Then we have
$\sum_{\begin{subarray}{c}s,t\in S\\ s\neq 0,t\neq 0\end{subarray}}\left<\Phi(\lfloor s^{-1}\rfloor\lfloor t\rfloor)h_{t},h_{s}\right$> $=\sum_{\begin{subarray}{c}s,t\in S\\ s\neq 0,t\neq 0\end{subarray}}\left<V^{*}\pi(\lfloor s^{-1}\rfloor)\pi(\lfloor t\rfloor)Vh_{t},h_{s}\right$>
$=\sum_{\begin{subarray}{c}s,t\in S\\ s\neq 0,t\neq 0\end{subarray}}\left<\pi(\lfloor t\rfloor)Vh_{t},\pi(\lfloor s^{-1}\rfloor)^{\dagger}Vh_{s}\right$>
$=\left<\sum_{\begin{subarray}{c}t\in S\\ t\neq 0\end{subarray}}\pi(\lfloor t\rfloor)Vh_{t},\sum_{\begin{subarray}{c}s\in S\\ s\neq 0\end{subarray}}\pi(\lfloor s\rfloor)Vh_{s}\right$> (95)
$\geq 0,$
where in Eq (95) we have used the fact that $\pi$ is a *-homomorphism. This shows that $\Phi$ is a positive definite map. This completes the proof. $\Box$
III.4 Choi theorem on Completely positive maps and Bochner’s theorem on positive definite maps on finite dimensional matrix algebra
As discussed earlier, the the matrix algebra $M_{n}(\mathbb{C})$ is a contracted algebra of the inverse semigroup $S_{M}=\{e_{ij}\}_{i,j=1}^{n}\cup\{0\}$. So, from Theorem (5) it is clear that a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ is positive definite if and only if it completely positive. Now from Bochner’s Theorem (4) we have that a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ is positive definite if and only if its Fourier transform $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(M_{m}(\mathbb{C}))$. From the discussions of Proposition (B), the inverse semigroup $S_{M}$ has two $\mathcal{D}$-classes, namely $D_{0}=\{0\}$ and $D_{1}=\{e_{ij}\}_{i,j=1}^{n}$ with $D_{1}$ being the relevant one to us. Let us fix the idempotent $e_{11}$. Now the maximal subgroup at $e_{11}$ is $G_{e_{11}}:=\{e_{11}\}$. Also we have $\lfloor e_{ij}\rfloor=e_{ij}$, $\mathrm{ran}(e_{ij})=e_{ii}$, $\mathrm{dom}(e_{ij})=e_{jj}$, $p_{\mathrm{ran}(e_{ij})}=e_{i1}$, $p_{\mathrm{dom}(e_{ij})}=e_{j1}$. Now from the definition of the induced representation $\overline{\rho}_{k}$, we have
$\overline{\rho}_{1}(e_{ij})$ $=E_{e_{ii},e_{jj}}\otimes\rho_{1}(e_{1i}e_{ij}e_{j1})$
$=E_{e_{ii},e_{jj}}\otimes\rho_{1}(e_{11}).$
Since $\rho_{1}$ is irreducible unitary representation of the singleton group $G_{e_{11}}=\{e_{11}\}$, we have $\rho_{1}(e_{11})=1$. Identifying $E_{e_{ii},e_{jj}}$ with $e_{ij}$, we have $\overline{\rho}_{1}(e_{ij})=e_{ij}$. From the definition of Fourier transform of $\Phi$ (see Eq. (41)),
$\widehat{\Phi}(\overline{\rho}_{1})$ $=\sum_{s\in S_{M},s\neq 0}\overline{\rho}_{1}(\lfloor s\rfloor)\otimes\widetilde{\Phi}(\lfloor s\rfloor)$
$=\sum_{i,j=1}^{n}\overline{\rho}_{1}(e_{ij})\otimes\Phi(e_{ij})$ (96)
$=\sum_{i,j=1}^{n}e_{ij}\otimes\Phi(e_{ij}),$ (97)
where in Eq. (96) we have used the fact that $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$. From Eq. (97) we see that $\widehat{\Phi}(\overline{\rho}_{1})$ is the Choi matrix of the linear map $\Phi$. Hence in the context of finite dimensional matrix algebra, Bochner’s theorem reduces to the Choi theorem on completely positive maps.
III.5 Relation between complete positivity of maps and positivity of their Fourier transforms
In the previous subsection, we have shown that the Bochner’s theorem on positive definite maps on matrix algebra reduces to the Choi theorem for completely positive(CP) maps. It should be noted that the Fourier transforms appeared in Bochner’s theorem are with respect to the induced representations arising from the irreducible unitary representations of the maximal subgroups of the finite inverse semigroup. So it interesting to know if the following statement is true: Let $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ be a completely positive map and let $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ be a representation. Then, is $\Phi$ completely positive if and only if $\widehat{\Phi}(\rho)$ is positive semidefinite? In the following theorem we provide a necessary and sufficient condition on the representation $\rho$ such that we have an affirmative answer to the question.
Theorem 6.
The CP vs. Positivity Correspondence between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier transform $\widehat{\Phi}(\rho)$ with respect to a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ holds if and only if $\rho$ is of the form $\rho(X)=UXU^{\dagger}$ for all $X\in M_{m}(\mathbb{C})$, where $U:\mathbb{C}^{m}\to\mathbb{C}^{d_{\rho}}$ is unitary with $d_{\rho}=m$.
Proof: From the definition, we have
$\widehat{\Phi}(\rho)$ $=\sum_{i,j}\rho(e_{ij})\otimes\Phi(e_{ij})$
$=(\mathrm{id}\otimes\Phi)\sum_{i,j}\rho(e_{ij})\otimes e_{ij}.$
Now it can be shown that $\sum_{i,j}\rho(e_{ij})\otimes e_{ij}=\tau(C_{\rho^{*}})$, where $\rho^{*}$ is the adjoint of the representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$, $C_{\rho^{*}}$ is the Choi matrix of $\rho^{*}$ and $\tau$ represents transposition operation. With this relation, the above equation becomes
$\widehat{\Phi}(\rho)$ $=(\mathrm{id}\otimes\Phi)\tau(C_{\rho^{*}})=(\tau\otimes\Phi\circ\tau)C_{\rho^{*}}$
$\iff\tau(\widehat{\Phi}(\rho))=(\mathrm{id}\otimes\tau\circ\Phi\circ\tau)C_{\rho^{*}}.$
As $\tau(\widehat{\Phi}(\rho))$ is positive semidefinite if and only if $\widehat{\Phi}(\rho)$ is positive semidefinite, and $\Phi$ is CP if and only if $\tau\circ\Phi\circ\tau$ is CP, then by using Theorem (1) of [21] (Also see [22]), we have that the CP vs positivity correspondence holds between $\Phi$ and $\widehat{\Phi}(\rho)$ if and only if $\rho^{*}$ is a complete order isomorphism if and only if $\rho$ is a complete order isomorphism. Since $\rho$ is a representation, it must be of the form $\rho(X)=UXU^{\dagger}$ for all $X\in M_{m}(\mathbb{C})$, where $U:\mathbb{C}^{m}\to\mathbb{C}^{d_{\rho}}$ is unitary with $d_{\rho}=m$. This completes the proof. $\Box$
IV Conclusion
The Choi-Jamiołkowski isomorphism is an essential tool in the study of various mathematical objects in quantum information theory. In this work, we have defined convolution product between maps from the contracted semigroup algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ to an arbitrary associative algebra $\mathcal{A}$. The convolution product makes the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to $\mathcal{A}$ an associative algebra. Then we have proved that the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ and the tensor product algebra $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ are isomorphic. In the specific case of matrix unit semigroup, our result shows that the product in the space of maps on the matrix algebras which is preserved by the Choi-Jamiołkowski isomorphism is convolution. Then we defined Fourier transform of a map from $\mathbb{C}_{0}[S]$ to $M_{n}(\mathbb{C})$ and derived the Fourier inversion formula when $S$ is a finite inverse semigroup. In the case of the inverse semigroup of matrix units, the Fourier transformation of a map with respect to the identity representation becomes the Choi matrix of the map and the Fourier inversion formula becomes the Choi inversion formula. We have also derived the Plancherel formula in the context of finite inverse semigroup. Then, by defining the notion of matrix valued positive definite maps, we have proved Bochner’s theorem in the context of finite inverse semigroup. Also, we have proved Schur’s orthogonality relation in the context of finite inverse semigroup which played a crucial role in the proof of Bochner’s theorem. Then we have demonstrated how Bochner’s theorem reduces to Choi theorem on completely positive maps when we consider the inverse semigroup of matrix units. Finally, we have derived the necessary and sufficient condition on a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$.
Appendix A Proof of Plancherel formula for finite groups
The Plancherel formula can be derived as follows:
$\sum_{g\in G}\Phi(g^{-1})\Psi(g)$ $=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}(\rho(g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]}\Psi(g)$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}(\rho(g)\otimes\mathbb{I})\widehat{\Phi}(\rho)(\mathbb{I}\otimes\Psi(g))\big{]}$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}\widehat{\Phi}(\rho)(\mathbb{I}\otimes\Psi(g))(\rho(g)\otimes\mathbb{I})\big{]}$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}\widehat{\Phi}(\rho)(\rho(g)\otimes\Psi(g))\big{]}$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}\widehat{\Phi}(\rho)\widehat{\Psi}(\rho)\big{]}.$
This completes the proof.
Appendix B Proof of Theorem (A)
[…]
$a\cdot a^{\prime}$ $=\bigg{(}\sum_{i}g_{i}\otimes\mathrm{\Phi}(g_{i})\bigg{)}\cdot\bigg{(}\sum_{j}g_{j}\otimes\mathrm{\Phi}^{\prime}(g_{j})\bigg{)}$
$=\sum_{i,j}g_{i}g_{j}\otimes\mathrm{\Phi}(g_{i})\mathrm{\Phi}^{\prime}(g_{j}).$ (98)
Now in the group multiplication table, along a row or column, each element of the group appears once. So it is guaranteed that the terms $g_{i}g_{j}$ for all possible values of $i$ and $j$ produces all the elements in the group. We collect together all the terms in the summation of Eq. (98) for which $g_{i}g_{j}$ is equal to a particular group element say $g_{k}$ and rewrite the summation as follows:
$a\cdot a^{\prime}$ $=\sum_{k}g_{k}\otimes\bigg{(}\sum_{\begin{subarray}{c}p,q:\\ g_{p}g_{q}=g_{k}\end{subarray}}\mathrm{\Phi}(g_{p})\mathrm{\Phi}^{\prime}(g_{q})\bigg{)},$ (99)
where the summation inside the big bracket is over those pair of elements $(g_{p},g_{q})$ such that $g_{p}g_{q}=g_{k}$. Note that the term inside the bracket is nothing but the convolution of the maps $\mathrm{\Phi}$ and $\mathrm{\Phi}^{\prime}$:
$\mathrm{\Phi}*\mathrm{\Phi}^{\prime}(g_{k})=\sum_{\begin{subarray}{c}p,q:\\ g_{p}g_{q}=g_{k}\end{subarray}}\mathrm{\Phi}(g_{p})\mathrm{\Phi}^{\prime}(g_{q}).$ (100)
Using Eq. (100) in Eq. (99), we finally have
$a\cdot a^{\prime}$ $=\sum_{k}g_{k}\otimes(\mathrm{\Phi}*\mathrm{\Phi}^{\prime})(g_{k}),$ (101)
which completes the proof.
Appendix C Proof of Proposition( 6)
Proposition 6.
For a finite inverse semigroup $S$ and $z\neq s,t\in S$, let $ss^{-1}=tt^{-1}$. Then $s^{-1}t$ is an idempotent if and only if $s=t$.
Proof: When $s=t$, trivially $s^{-1}t$ is an idempotent.
Now, let $s^{-1}t$ is an idempotent. Then
$s^{-1}ts^{-1}t=s^{-1}t$ (102)
$s^{-1}t=t^{-1}s$ (103)
$(s^{-1}t)^{-1}(s^{-1}t)=(s^{-1}t)(s^{-1}t)^{-1}.$ (104)
Using Eq. (104) and the assumption $ss^{-1}=tt^{-1}$, we have
$s^{-1}s=t^{-1}t.$ (105)
Now we use Eq. (103) in Eq. (102) to obtain
$s^{-1}tt^{-1}s=s^{-1}t$
$\iff s^{-1}ss^{-1}s=s^{-1}t$
$\iff s^{-1}s=s^{-1}t$
$\iff ts^{-1}s=ts^{-1}t$
$\iff tt^{-1}t=ts^{-1}t$ (106)
$\iff t=ts^{-1}t,$ (107)
where in Eq. (106) we have used eq. (105). Since in a finite inverse semigroup every element has a unique inverse, we conclude from Eq. (107) that $s=t$. This completes the proof. $\Box$
Appendix D Proof of Bochner’s theorem for matrix valued maps on finite group
We recall that the inversion formula [see Eq. (33)] for maps on finite group is given by
$\Phi(g)=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{-1})\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]},$
$\Phi(g^{-1}g^{\prime})=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{\prime-1}g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]}.$
Now, using the above equation we have
$\sum_{g,g^{\prime}\in G}\overline{\left<\Phi(g^{-1}g^{\prime})h_{g^{\prime}},h_{g}\right>}$ $=\sum_{g,g^{\prime}\in G}\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\overline{\left<\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{\prime-1}g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]}h_{g^{\prime}},h_{g}\right>}$
$=\sum_{g,g^{\prime}\in G}\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\left<h_{g},\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{\prime-1}g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]}h_{g^{\prime}}\right$>
$=\sum_{g,g^{\prime}\in G}\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\mathrm{tr}\left(\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{\prime-1}g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]}h_{g^{\prime}}h_{g}^{\dagger}\right)$
$=\sum_{g,g^{\prime}\in G}\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\mathrm{tr}\left((\rho(g^{\prime})^{\dagger}\rho(g))\otimes h_{g^{\prime}}h_{g}^{\dagger})\widehat{\Phi}(\rho)\right)$
$=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\mathrm{tr}\left(\left(\sum_{g^{\prime}\in G}\rho(g^{\prime})^{\dagger}\otimes h_{g^{\prime}}\right)\left(\sum_{g\in G}\rho(g)\otimes h_{g}^{\dagger}\right)\widehat{\Phi}(\rho)\right)$
$=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\mathrm{tr}\left(A(\rho,h)^{\dagger}A(\rho,h)\widehat{\Phi}(\rho)\right),$ (109)
where $A(\rho,h):=\sum_{g\in G}\rho(g)\otimes h_{g}^{\dagger}$ and $h:=\{h_{g}\}_{g\in G}$ Now, if we assume that $\widehat{\Phi}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU}(G)$, then we have $\sum_{g,g^{\prime}\in G}\overline{\left<\Phi(g^{-1}g^{\prime})h_{g^{\prime}},h_{g}\right>}\geq 0$ for all subset $h\subset\mathbb{C}^{n}$ implying that $\Phi$ is a positive definite map. This completes the proof in one direction.
For the other direction, let us fix $\rho^{0}\in\mathrm{IrrU}(G)$ and choose $h^{0}$ as the subset consisting of the vectors $h^{0}_{g}$ of the form
$h^{0}_{g}=\operatorname{tr}_{d_{\rho^{0}}}\left(\rho^{0}(g)\otimes\mathbb{I})B^{\dagger}\right),$ (110)
where $B\in M_{d_{\rho^{0}}}(\mathbb{C})\otimes\mathbb{C}^{n}$ and $\operatorname{tr}_{d_{\rho^{0}}}$ represents trace over the space $M_{d_{\rho^{0}}}(\mathbb{C})$. Now, using Schur orthogonality relations, we have
$A(\rho,h^{0})$ $=\begin{cases}\frac{|G|}{d_{\rho^{0}}}B&\text{when}~\rho=\rho^{0},\\ 0&\text{when}~\rho\neq\rho^{0}.\end{cases}$ (111)
Now by substituting the above form of $A(\rho,h^{0})$ in Eq. (109), we get
$\sum_{g,g^{\prime}\in G}\overline{\left<\Phi(g^{-1}g^{\prime})h^{0}_{g^{\prime}},h^{0}_{g}\right>}$ $=\frac{d_{\rho^{0}}}{|G|}\mathrm{tr}\left(A(\rho^{0},h^{0})^{\dagger}A(\rho^{0},h^{0})\widehat{\Phi}(\rho)\right)$ (112)
$=\frac{|G|}{d_{\rho^{0}}}\mathrm{tr}\left(B^{\dagger}B\widehat{\Phi}(\rho)\right).$ (113)
Now if we assume that $\Phi:G\to M_{n}(\mathbb{C})$ is positive definite then the left hand side of the above equation is non-negative. Since, $B$ and $\rho^{0}$ are chosen arbitrarily, $\widehat{\Phi}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU}(G)$. This completes the proof.